### v2

# Long-term impact of the magnetic-field strength on the evolution and electromagnetic emission by neutron-star merger remnants

Abstract
Numerical simulations are essential to understand the complex physics accompanying the merger of binary systems of neutron stars. However, these simulations become computationally challenging when they have to model the merger remnants on timescales over which secular phenomena, such as the launching of magnetically driven outflows, develop. To tackle these challenges, we have recently developed a hybrid approach that combines, via a hand-off transition, a fully general-relativistic code (FIL) with a more efficient code making use of the conformally flat approximation (BHAC+). We here report important additional developments of BHAC+ consisting of the inclusion of gravitational-wave radiation-reaction contributions and of higher-order formulations of the equations of general-relativistic magnetohydrodynamics. Both improvements have allowed us to explore BNS merger remnants with high accuracy and over timescales that would have been computationally prohibitive otherwise. More specifically, we have investigated the impact of the magnetic-field strength on the long-term (i.e., $\sim 200\,{\rm ms}$) and high-resolution (i.e., $150\,{\rm m}$) evolutions of the “magnetar” resulting from the merger of two neutron stars with a realistic equation of state. In this way, and for sufficiently large magnetic fields, we observe the weakening or suppression of differential rotation and the generation of magnetic flares in the outer layers of the remnant. These flares, driven mostly by the Parker instability, are responsible for intense and collimated Poynting flux outbursts and mass ejections. This novel phenomenology offers the possibility of seeking corresponding signatures from the observations of short gamma-ray bursts and hence revealing the existence of a long-lived strongly magnetized remnant.
I Introduction
The observation of the gravitational-wave (GW) event GW170817 and the corresponding electromagnetic (EM) emission The LIGO Scientific Collaboration and The Virgo Collaboration (2017); Abbott et al. (2017); The LIGO Scientific Collaboration et al. (2017) marked the dawn of a new era in multi-messenger astronomy, combining GW and EM observations. This breakthrough in multi-messenger signals provides us with unprecedented opportunities to address long-standing questions in physics, such as the behavior of the equation of state (EOS) for extremely dense nuclear matter, the mechanisms behind the launching of relativistic jets and short gamma-ray bursts (GRBs) Rezzolla et al. (2011); Just et al. (2016); Ciolfi (2020); Hayashi et al. (2022); Baiotti and Rezzolla (2017); Paschalidis (2017); Murguia-Berthier et al. (2016); Lazzati et al. (2018), and the origin of heavy elements in the universe Metzger et al. (2010); Bovard et al. (2017); Smartt and Chen (2017); Papenfort et al. (2018); Combi and Siegel (2023); Fujibayashi et al. (2023); Kawaguchi et al. (2023). To answer these questions thoroughly, it is necessary to model the highly nonlinear processes involved in these multi-messenger events with accuracy and realism. This requires self-consistent numerical simulations that solve the Einstein equations, general-relativistic magnetohydrodynamics (GRMHD), and neutrino radiation transport, while incorporating realistic tabulated EOSs. These advanced techniques are vital for capturing the intricate details of binary neutron-star (BNS) merger dynamics and bridging theoretical predictions with observational data.
Although significant progress has been achieved in numerically modelling BNS mergers in recent years, two major aspects still require further investigation. The first pertains to high-resolution simulations that focus on MHD processes, such as the Kelvin-Helmholtz instability (KHI) Rasio and Shapiro (1999); Price and Rosswog (2006); Anderson et al. (2008); Baiotti et al. (2008); Chabanov et al. (2023), the $\alpha$-$\Omega$ dynamo Ruediger and Kichatinov (1993); Bonanno et al. (2003); Most (2023); Musolino et al. (2024a), the magnetorotational instability (MRI) Velikhov (1959); Chandrasekhar (1960, 1960); Siegel et al. (2013), and magnetic winding/braking Siegel et al. (2014); Kiuchi et al. (2018, 2023), which may all occur during BNS mergers. The impact of these MHD processes on the evolution of BNS mergers is enormous, but hard to quantify as several of them are highly sensitive to grid resolution. For instance, the MRI contributes to the exponential growth of large-scale poloidal magnetic fields, facilitating angular momentum transfer, influencing black-hole collapse time, driving strong postmerger winds, and potentially affecting relativistic jet formation. However, the MRI’s effectiveness depends sensitively on simulation resolution and on the magnetic-field topology generated by the KHI process, which is also highly resolution dependent. Given the critical dependence on resolution for both instability processes, and the lack of convergence in current simulations, there is an urgent need for high-resolution simulations. Yet, these simulations are computationally expensive and time-intensive, necessitating efficient approaches to reach feasible solutions.
The second aspect involves an urgent need for long-term, large-scale simulations to consistently capture all multi-messenger signals produced in compact binary merger events, e.g., GWs Abbott et al. (2017, 2019), GRB multi-band EM emission Fermi-LAT Collaboration (2017); Goldstein et al. (2017), multi-band afterglow emission Hallinan et al. (2017); Troja et al. (2017); Margutti et al. (2018); Mooley et al. (2018); Ghirlanda et al. (2019); Musolino et al. (2024b), and kilonova observations Chornock and et al. (2017); Li and Paczyński (1998); Drout et al. (2017); Nicholl et al. (2017); Pian et al. (2017); Soares-Santos et al. (2017). These general-relativistic simulations must cover timescales of seconds to account for the delay from the GW to the associated GRB signal, which is $1.7\,\mathrm{s}$ in the case of GW170817 Goldstein et al. (2017), and hints at a collapse time of about $1.0\,\mathrm{s}$ Gill et al. (2019); Murguia-Berthier et al. (2021); Nathanail et al. (2021a). Even longer timescales, although not necessarily in full general relativity (GR), must be span to describe the afterglow signal appearing days later Margutti et al. (2017) and lasting months Troja et al. (2019). This presents challenges in maintaining the same simulation method over extended periods. Hence, transitions between codes (hand-off) at different stages offer a versatile approach, enabling various physical processes to be modelled while leveraging the unique strengths of different codes (see, e.g.,  Lopez Armengol et al. (2022); Gottlieb et al. (2023a, b, 2023); Ng et al. (2024a); Ennoggi et al. (2025)).
Using the Conformal Flatness Condition (CFC) to solve the elliptic sector of the Einstein equations offers an efficient approach to these challenges, as it reduces the need for frequent spacetime computations when the spacetime is not highly dynamic Ng et al. (2024a). The CFC approximation has been successfully applied in simulations of core-collapse supernovae Dimmelmeier et al. (2002a); Ott et al. (2007); Müller (2015); Cheong et al. (2023); Ng et al. (2024b), rapidly rotating neutron stars Dimmelmeier et al. (2002b); Cordero-Carrión et al. (2009); Ng et al. (2021); Cheong et al. (2024), and BNS mergers Bauswein et al. (2012, 2021); Lioutas et al. (2024). Our previous work Ng et al. (2024a), further demonstrated that the extended-CFC (xCFC) scheme Cordero-Carrión et al. (2009) implemented in our new , multi-coordinate and multi-dimensional GRMHD code BHAC+ can effectively handle the output of another code, FIL Most et al. (2019, 2021); Chabanov et al. (2023); Musolino and Rezzolla (2024); Chabanov and Rezzolla (2023); Ng et al. (2024), which solves the Einstein equations using full general relativity. Specifically, we have shown that the CFC approximation provides nearly identical solutions in the head-on collision of two neutron stars, not only in the evolution of maximum rest-mass density, but also in the two-dimensional (2D) rest-mass density and temperature distributions, when compared to the full-GR code FIL.
In this work, we present the implementation, testing and use of three major extensions to BHAC+. First, we go beyond the limitations of CFC by introducing a phenomenological GW radiation-reaction (RR) scheme into BHAC+ Ng et al. (2024a). This addition compensates for the effects of residual GW emission and enables more accurate modelling of the fluid and magnetic-field dynamics during the postmerger phase. Our calculations of the RR corrections are based on Refs. Faye and Schäfer (2003); Oechslin et al. (2007), with significant improvements to both the accuracy of the scheme and its coupling to the xCFC framework. Second, we have improved several numerical algorithms to higher-order ones, such as replacing the second-order finite-volume method with a fourth-order finite-difference approach, and switching from the Piecewise Parabolic Method (PPM) reconstruction Colella and Woodward (1984) to a WENO-Z+ reconstruction Acker et al. (2016). Third, the hand-off method has been extended for parallel execution and optimised significantly by reducing the memory requirements.
Leveraging on these improvements, and exploiting the results of simulations of the inspiral and merger of BNSs computed by the full-GR code FIL, we have conducted with BHAC+ long-term and three-dimensional simulations of the merger remnant, starting from different postmerger stages to verify the consistency between the results of FIL and BHAC+. We found that the xCFC scheme remains highly accurate, robust, and efficient in capturing the long-term postmerger evolution of the fluid and magnetic fields. This holds true despite the differences in numerical methods for evolving spacetime and handling the divergence-free condition between BHAC+ and FIL.
Using this approach, we were able to perform four long-term simulations of neutron-star merger remnants differentiating either in the strength of the magnetic field (high/low) and in the resolution ($300/150\,{\rm m}$). These simulations have allowed us to assess the impact of the magnetic-field strength on the evolution and EM emission by neutron-star merger remnants, revealing, for instance, that very strong magnetic fields can moderate or suppress the differential rotation in the remnant. Additionally, they can lead to the generation of a strong and collimated Poynting flux after the Parker instability criterion is met and to the generation of magnetic flares in the outer layers of the remnant that could accompany the phenomenology of short GRBs.
The paper is organised as follows. In Sec. II, we describe the mathematical formulation of the CFC approximation and outline the coupling of a novel RR module to the xCFC scheme. The basic numerical setup and the details of the hand-off are presented in Sec. III, while in Sec. IV, we validate the hand-off and RR through a series of simulations. We focus on the long-term impact of strong magnetic fields on the merger remnant in Sec. V. Finally, we discuss our results in Sec. VI and suggest possible directions for future research.
Throughout this paper, unless stated otherwise, we adopt units in which $c=G=M_{\odot}=k_{\rm B}=\epsilon_{0}=\mu_{0}=1$, except for the coordinates. Latin indices denote spatial components (from $1$ to $3$), while Greek indices indicate spacetime components (from 00 to $3$).
II Mathematical Setup
II.1 The CFC approximation and extended CFC scheme
We recall that the CFC reduces the conformally-related spatial metric $\tilde{\gamma}_{ij}$ to the flat metric $f_{ij}$. Thus, the spatial metric can be expressed in terms of the conformal factor $\psi$ and the flat metric, i.e.,
$\gamma_{ij}=\psi^{4}\tilde{\gamma}_{ij}\,.$ (1)
This approximation suppresses the radiative degrees of freedom in the Einstein equations and is thus known as the “waveless approximation”. By employing the maximal slicing condition $K=\gamma^{ij}K_{ij}=0$, where $K_{ij}$ is the extrinsic curvature, the Hamiltonian and momentum-constraint equations reduce to a set of coupled nonlinear elliptic differential equations. This set of elliptic equations can be further refined into a well-established scheme known as the xCFC scheme Cordero-Carrión et al. (2009).
In the context of the xCFC scheme implemented in BHAC+ Ng et al. (2024a), the resulting set of elliptic equations is given by
$\hat{\Delta}X^{i}+\frac{1}{3}\hat{\nabla}^{i}\left(\hat{\nabla}_{j}X^{j}\right)=8\pi f^{ij}\tilde{S}_{j}\,,$ (2)
$\hat{\Delta}\psi=-2\pi\psi^{-1}\tilde{E}-\frac{1}{8}\psi^{-7}f_{ik}f_{jl}\hat{A}^{kl}\hat{A}^{ij}\,,$ (3)
$\hat{\Delta}(\alpha\psi)=(\alpha\psi)\left[2\pi\psi^{-2}(\tilde{E}+2\tilde{S})+\frac{7}{8}\psi^{-8}f_{ik}f_{jl}\hat{A}^{kl}\hat{A}^{ij}\right]\,,$ (4)
$\hat{\Delta}\beta^{i}+\frac{1}{3}\hat{\nabla}^{i}\left(\hat{\nabla}_{j}\beta^{j}\right)=16\pi\alpha\psi^{-6}f^{ij}\tilde{S}_{j}+2\hat{A}^{ij}\hat{\nabla}_{j}\left(\psi^{-6}\alpha\right)\,,$ (5)
where $\tilde{S}_{j}:=\psi^{6}S_{j}$, $\tilde{E}:=\psi^{6}E$, and $S$ and $E$ are the conserved momentum flux and the conserved energy density, respectively. The tensor $\hat{A}^{ij}$ is the conformal extrinsic curvature, given by
$\hat{A}^{ij}\approx\hat{\nabla}^{i}X^{j}+\hat{\nabla}^{j}X^{i}-\frac{2}{3}\hat{\nabla}_{k}X^{k}f^{ij}\,.$ (6)
(for more details on the mathematical formulation and technical aspects, we refer the reader to Refs. Cordero-Carrión et al. (2009); Cheong et al. (2020, 2021); Ng et al. (2024a)).
A crucial aspect of the CFC approach that makes it computationally attractive is that the xCFC equations are not necessarily updated at every time-step or sub-step of the Runge-Kutta method. Instead, we can solve this set of elliptic equations at much lower frequency, which we quantify with the ratio $\chi:=\Delta t_{\rm met}/\Delta t_{{}_{\rm MHD}}$, where $\Delta t_{\rm met}$ is the time interval between two metric updates, and $\Delta t_{{}_{\rm MHD}}$ is the time-step of the GRMHD simulation. The value of $\chi$ depends on the resolution of the simulation, the timescale of spacetime changes, and the degree of conformal flatness, but can easily reach values $\chi\simeq 10$ and even be as large as $\chi\simeq 40$ at later stages in the postmerger evolution (see Appendix C for a discussion on the impact of different values of $\chi$ and the postmerger evolution). The actual solution of the xCFC equations is made employing an efficient and low-memory cell-centered multigrid solver Cheong et al. (2020, 2021), as a result, the impact of the solution of the elliptic sector is only $\sim 40\%$ of the total cost of a reference time-step (similarly, the solution of the full Einstein equations represents $\sim 40\%$ of the time-step from FIL). In addition, when solving the set of GRMHD equations, the time-step size needs not be CFL-limited by the speed of light, but by the relevant MHD speeds, which are $\sim 30\%$ smaller. As a result, speed-ups of the order of $3-4$ are seen in our hybrid approach (see also Appendix C and Tab. 3 for a detailed discussion).
II.2 GW radiation-reaction in the xCFC scheme
Some of the initial work on the inclusion of RR on the energy and momentum budget in numerical codes solving the post-Newtonian (PN) formulation of the hydrodynamic equations was reported in Refs. Blanchet et al. (1990); Rezzolla et al. (1999); Faye and Schäfer (2003). These approaches were then coupled phenomenologically with the CFC scheme thereby compensating for the waveless property of the CFC approximation Oechslin et al. (2007); Lioutas et al. (2024). By implementing the 3.5 PN order formalism, but including only all 2.5PN terms and 3.5 PN corrections related to the gravitational potential and its derivatives in the metric components, the RR scheme provides a gravitational description with modified metric components coupled to hydrodynamics to approximate the system’s energy-momentum loss via GWs.
However, the previous approaches have several disadvantages and require various approximations (see, e.g., Ref Oechslin et al. (2007)). First of all, loss of accuracy can result if the time-step is not sufficiently small. This is the result of the need to perform two additional time-derivatives of the first time-derivative of the quadrupole moment $I^{[1]}_{ij}$ which are computed using finite-differencing as outlined in Ref. Oechslin et al. (2007) (hereafter, $X^{[k]}$ denotes the $k$-th order time-derivative of $X$ with respect to the coordinate time). To maintain accuracy and stability, particularly when there are rapid changes in the quadrupole moment $I_{ij}$, time-derivatives must be updated at every sub-step of the time integration and with a very small time-step. Secondly, because GW emission remains significant at $\sim 10-20\,\rm{ms}$ after merger (see also Fig. 1), high-order time derivatives of the Newtonian mass quadrupole are not small and need to be properly accounted for. Finally, previous work (e.g., Refs. Oechslin et al. (2007); Lioutas et al. (2024)) has not provided a close comparison between a CFC scheme including RR and a full-GR scheme, making it hard to judge the validity of some of the assumptions made in implementing the RR terms. To the best of our knowledge, this is the first time that a detailed comparison is presented on the role of the RR terms in the CFC approximation. Given the intricacies and subtleties in computing these terms and those discussed above, we hope that this information will be of use for future studies employing the CFC approximation (and the RR corrections) in BNS postmerger simulations.
In view of these considerations, we have taken a fresh new look at the whole approach of including RR terms within the 3.5 PN order formalism introduced in Ref. Oechslin et al. (2007), and hence derived a modified RR scheme. In this approach, we employ terms based on Ref. Faye and Schäfer (2003), but include the magnetic energy in the estimation of the matter sources, and couple the scheme to the metric components of the xCFC equations. We therefore solve the following set of nonlinear elliptic equations in terms of the six scalar potentials $\mathcal{U}_{*},\mathcal{U}_{*i},\mathcal{R}$, and $\mathcal{R}_{2}$ (we here consider $c=1$ but we report it explicitly in Appendix B to keep track of the various PN orders)
$\Delta\mathcal{U}_{*}=-4\pi\sigma\,,$ (7a)
$\Delta\mathcal{U}_{*i}=-4\pi\sigma w_{i}\,,$ (7b)
$\Delta\mathcal{R}=-4\pi Q_{ij}^{[3]}x^{i}\partial_{j}\sigma\,,$ (7c)
$\Delta\mathcal{R}_{2}=-4\pi\sigma\left(Q_{ij}^{[3]}x^{i}\partial_{j}\mathcal{U}_{*}-3Q_{ij}^{[3]}w_{i}w_{j}-\mathcal{R}\right)\,,$ (7d)
where $\sigma:=T^{ii}+T^{00}$ is the 1 PN + 3.5 PN mass density to replace the conserved rest-mass density $D^{*}:=D\psi^{6}=\rho W\psi^{6}$ Oechslin et al. (2007) (see Eq. (28) in Appendix B), given that $T^{\mu\nu}$ is the energy-momentum tensor, and $w_{i}:=hu_{i}$ is the momentum per unit rest-mass in the fluid frame as defined in Eq. (4.2) of Blanchet et al. (1990). Furthermore, $u_{\mu}$ is the covariant fluid four-velocity, and $h(\rho,T,Y_{p})=1+\epsilon+(p+b^{2})/\rho$ is the total relativistic specific enthalpy, thus including the contribution from EM fields ($b^{2}=b^{\mu}b_{\mu}$ is the square of the magnetic field strength in the fluid frame) in addition to that from the specific internal energy $\epsilon$, the fluid pressure $p$, rest-mass density $\rho$ and proton fraction $Y_{p}$.
Equations (7a)-(7d) correspond to Eqs. (4.31j), (4.31k), (4.31n), and (4.31cc) in Faye and Schäfer (2003), where we keep all 2.5 PN quantities and 3.5 PN corrections related to the PN potential $\mathcal{R}_{2}$ and its derivatives, which depends on the quantity $Q_{ij}^{[3]}$ [see Eq. (8)]. Note also that Eq. (7d) includes the term $\sim 3Q_{ij}^{[3]}w_{i}w_{j}$ that was omitted in Ref. Oechslin et al. (2007) but that actually leads to significant improvements in the accuracy of the RR scheme [see Appendix B and the derivation of Eq. (32)]. The quantity $Q_{ij}^{[3]}$ can be shown to be related to the third time-derivative of the Newtonian mass quadrupole (we will ignore here mass-current quadrupole but see Ref. Rezzolla et al. (1999) for how to include these corrections), where the two are related as Blanchet et al. (1990)
$Q_{ij}^{[3]}=I_{ij}^{[3]}+\mathcal{O}\left(\frac{1}{c^{2}}\right)\,.$ (8)
Note that while $I_{ij}^{[3]}$ is a purely Newtonian (i.e., 0 PN) quantity, $Q_{ij}^{[3]}$ is a genuine PN quantity entering at 2.5 PN in our scheme and computed as [see Eq. (5.13) in Blanchet et al. (1990) and Eq. (B4) in Faye and Schäfer (2003)]
$Q_{ij}^{[3]}=$ $\Biggl{[}\int_{\mathcal{V}}\sqrt{\hat{\gamma}}d^{3}x\,D^{*}$ (9)
$2\left[w_{i}\partial_{j}\mathcal{U}_{*}-2w_{i}\left[\partial_{j}(h-1)-T\partial_{j}s\right]\right.$
$+\left.x^{i}w_{k}\partial_{k}\partial_{j}\mathcal{U}_{*}-x^{i}\partial_{k}\partial_{j}\mathcal{U}_{*k}\right]\Biggr{]}^{\mathrm{STF}}\,,$
where $\mathcal{V}$ is the numerical domain, $s$ is the specific entropy (entropy per baryon), $W$ the Lorentz factor, $\sqrt{\hat{\gamma}}$ the determinant of the spatial part of the flat metric, and the superscript ${\rm STF}$ refers to a symmetric trace-free tensor, i.e., $\left[A^{ij}\right]^{\rm STF}:=\frac{1}{2}(A^{ij}+A^{ji})-\frac{1}{3}\delta^{ij}A^{kk}$ denotes the symmetric trace-free part of the tensor.
We note that the use of $Q_{ij}^{[3]}$ computed as in Eq. (9) avoids numerical time derivatives altogether and should be preferred to the calculation of $I_{ij}^{[3]}$ from the actual second-order time-derivative of the analytically calculated expression for $I^{[1]}_{ij}$ proposed in Refs. Oechslin et al. (2007); Lioutas et al. (2024) and computed as in Ref. Oechslin et al. (2007)
$I^{[3]}_{ij}(t)=$ $2\left[I^{[1]}_{ij}(t+dt_{1})-[1+dt_{1}/dt_{2}]I^{[1]}_{ij}(t)+\right.$ (10)
$\left.[dt_{1}/dt_{2}]I^{[1]}_{ij}(t-dt_{2})\right]/\left[dt_{1}(dt_{1}+dt_{2})\right]\,,$
where $dt_{1}$ and $dt_{2}$ are the current and last time-steps, respectively (we note that the calculation of $I^{[1]}_{ij}$ itself requires time-derivatives of several quantities, e.g., $\rho$, $\psi$ and $\epsilon$). By computing $Q_{ij}^{[3]}$ instead of $I_{ij}^{[3]}$, it is possible to remove the source of potential numerical errors and the numerous (and sometimes arbitrary) approximations made in their computation, i.e., the neglect of several terms in the total time-derivatives of $\mathcal{U}$, $\partial_{i}\mathcal{U}$, $V^{i}$ and $h$ in the expression of $I_{ij}$, where $\mathcal{U}$ and $V^{i}$ are the Newtonian potential and the transport velocity, respectively Oechslin et al. (2007); Lioutas et al. (2024) [^1] We note that some of the approximations in Refs. Oechslin et al. (2007); Lioutas et al. (2024) have been made assuming that the neutron-star binaries are in the unrealistic condition of corotation. However, these assumptions may be inaccurate in the case of irrotational or spinning BNSs.. In addition, our experience has shown that the calculation of $I_{ij}^{[3]}$ is extremely inaccurate near discontinuities or sharp gradients (e.g., the stellar surface). Reducing drastically the time-step can improve the calculation of $I_{ij}^{[3]}$ but obviously increases considerably the computational costs, which we are instead interested in limiting. By contrast, computing $Q_{ij}^{[3]}$ is not only more accurate but also allows for a time-step that is large and equal to that of the xCFC solver, i.e., $\Delta t_{\rm met}=0.5\,M_{\odot}$, chosen for postmerger phase.
By following Eqs. (4.31a), (4.31b) in Ref. Faye and Schäfer (2003) and keeping only the terms involving $Q_{ij}^{[3]}$, the metric component $g_{00}$ for the CFC system gains a new RR term given by
$g_{00,{{}_{\rm RR}}}=-\frac{4}{5}\left(1-2\mathcal{U}_{*}\right)\left(Q_{ij}^{[3]}x^{i}\partial_{j}\mathcal{U}_{*}-\mathcal{R}\right)+\frac{4}{5}\mathcal{R}_{2}\,.$ (11)
Note that the form of $g_{00,{{}_{\rm RR}}}$ is different from the corresponding Eq. (A.37) in Oechslin et al. (2007) [or Eq. (A.5) in Ref. Lioutas et al. (2024)], which contains a typo Oechslin (we refer the reader to the derivation presented in Eqs. (30)–(33) in Appendix B for further details).
As a result, the new (primed) metric components modified by the inclusion of RR take the form
$g_{00}^{\prime}$ $:=g_{00}+g_{00,{{}_{\rm RR}}}\,,$ (12)
so that the new lapse function can be expressed as
$\alpha^{\prime}$ $=\sqrt{\alpha^{2}-g_{00,{{}_{\rm RR}}}+\beta^{\prime i}\beta^{\prime}_{i}-\beta^{i}\beta_{i}}\,,$ (13)
$\approx\sqrt{\alpha^{2}-g_{00,{{}_{\rm RR}}}}\,,$ (14)
where the second equality is obtained when assuming that $\beta^{\prime i}\beta^{\prime}_{i}\simeq\beta^{i}\beta_{i}$, which is reasonable since the leading order of $g_{0i,{{}_{\rm RR}}}$ is a 3.5 PN quantity and $|\beta^{\prime}_{i}-\beta_{i}|\ll 1$ in the postmerger phase.
In addition, we should remark that while there are RR corrections also in the spatial part of the metric, i.e., $g_{ij,{{}_{\rm RR}}}=\gamma_{ij,{{}_{\rm RR}}}=-(4/5)Q_{ij}^{[3]}$ [see Eqs. (4.31d) and (4.31m) in Ref. Faye and Schäfer (2003)] these terms are difficult to incorporate in our CFC approximation scheme. This is because of the intrinsically diagonal form of the CFC equations, which allows us to replace $\psi$ with the RR-corrected $\psi^{\prime}$, but does not give us access to the full three-metric $\gamma_{ij}^{\prime}$. Using only the diagonal terms in $\gamma_{ij,{{}_{\rm RR}}}$ is obviously mathematically inconsistent and, unsurprisingly, when we have tried to apply $\psi^{\prime}=:(\det{\gamma_{ij}^{\prime}})^{1/12}$ we have encountered significant numerical problems such as non-preserved divergence-free condition and less accurate conservation of rest-mass and energy-momentum. Hence, in our implementation of the RR terms, we simply consider the components of $\gamma_{ij,{{}_{\rm RR}}}$ as high-order PN terms and set them to zero. To the best of our knowledge Oechslin , these terms are neglected also in other implementations of the CFC scheme (e.g., Oechslin et al. (2007)), although they can and are employed to compute the GW signal.
Following Refs. Blanchet et al. (1990); Faye and Schäfer (2003), our coupling of the RR terms to the xCFC scheme is made using a three-dimensional (3D) Cartesian coordinate system in which we compute the RR equations after every update of the xCFC scheme. In particular, using a multigrid algorithm Cheong et al. (2021); Ng et al. (2024a); Lioutas et al. (2024), we first solve the two elliptic equations (7a) and (7b) from which we compute $Q_{ij}^{[3]}$ as in Eq. (9). Next, we solve Eqs. (7c), (7d), and (11) using the newly computed value for $Q_{ij}^{[3]}$. The full set of elliptic equations (including Eqs. (2)-(5)) is then solved iteratively until the maximum absolute value of the residual falls below a tolerance of $10^{-6}$, which is sufficient for convergence. A Robin boundary condition is chosen for this procedure.
Once the correction $g_{00,{{}_{\rm RR}}}$ is computed from Eq. (11), we obtain a modified lapse function $\alpha^{\prime}$, which replaces $\alpha$ in the matter solver of the GRMHD equations. However, when solving the xCFC metric equations, we continue to use the original lapse function $\alpha$ instead. This approach ensures that the RR impacts only the matter evolution and not the field variables. We have found that, at least for the tests presented here, not updating the lapse function for the xCFC equations does lead to a better match with the results of the full-GR code FIL. Additionally, we also keep the conformal factor $\psi$ unchanged, thus ensuring that metric modifications do not alter the conserved quantities on a given spatial hypersurface. Not doing so would introduce modifications of the three-metric $\gamma_{ij}$, thus compromising rest-mass conservation but also the divergence-free condition for the magnetic field. Finally, because Eq. (11) does not guarantee positivity, it is possible, under extreme conditions, that negative values of $g_{00,{{}_{\rm RR}}}$ appear in a small number of grid-points; in these cases we simply set $g_{00,{{}_{\rm RR}}}=0$.
III Numerical Setup
III.1 FIL, FUKA and BHAC+
We here briefly recall the basic elements of our computational infrastructure and report the relevant references for additional details.
FIL Most et al. (2019, 2021); Chabanov et al. (2023); Musolino and Rezzolla (2024); Chabanov and Rezzolla (2023); Ng et al. (2024) solves the hyperbolic sector of the Einstein equations Alcubierre (2006); Rezzolla and Zanotti (2013) using full GR schemes, including BSSNOK Shibata and Nakamura (1995); Baumgarte and Shapiro (1999), BSSNOK-Z4c Bernuzzi and Hilditch (2010), and CCZ4 Alic et al. (2012, 2013). Coupled with the EinsteinToolkit Loeffler et al. (2012); Zlochower and et al. (2022), it utilises the Carpet box-in-box Adaptive Mesh Refinement (AMR) driver in Cartesian coordinates Schnetter et al. (2006). Besides, it uses a fourth-order accurate conservative finite-difference High-Resolution Shock-Capturing (HRSC) scheme with WENO-Z reconstruction Del Zanna et al. (2007), an HLL Riemann solver, and a vector-potential-based magnetic evolving method Etienne et al. (2015).
The initial data of FIL is generated using the publicly available FUKA code Grandclement (2010); Papenfort et al. (2021); Tootle (2023), which computes the initial data timeslice by solving the eXtended Conformal Thin Sandwich system of equations Pfeiffer and York (2005); Papenfort et al. (2021).
BHAC+ Ng et al. (2024a) (which is an extension of the publicly available BHAC Porth et al. (2017); Olivares et al. (2019); Ripperda et al. (2019); Mpisketzis et al. (2024, 2024)) solves the elliptic sector of the Einstein equations using the xCFC scheme Cordero-Carrión et al. (2009), with a cell-centered multigrid solver Cheong et al. (2020, 2021) and block-based quadtree-octree AMR Xia et al. (2018); Keppens et al. (2021). For consistency with FIL, BHAC+ is extended to incorporate fourth-order finite-difference HRSC, the same reconstruction method, and the same Riemann solver. However, it employs a second-order convergent scheme for the xCFC metric solver (with or without RR) and a magnetic-field-based upwind constrained-transport method Olivares et al. (2019).
In the default setup, BHAC+ makes use of the WENO-Z reconstruction and fourth-order finite differencing and this leads to the closest matches with the results from FIL in terms of the evolution of maximum density, rotational profiles, and total EM energy. It is important to mention that both approaches to the violation of the divergence-free condition, namely, the use of a vector potential or of the constrained-transport method Evans and Hawley (1988), guarantee a magnetic field that is divergence free at machine precision. It is also worth noting that we have resolved an artefact reported in our previous work Ng et al. (2024a), where we observed that when performing an azimuthally averaged hand-off, the postmerger remnant experienced fluctuating and significantly lower temperatures in the core region. As discussed in Ng et al. (2024a), this behaviour was partly due to slightly inaccurate values of the specific internal energy $\epsilon$ generated by metric initialisation with the azimuthally averaged hand-off, but also to the poor resolution of the tabulated EOS, which has a strong dependence of the temperature $T$ on the specific internal energy $\epsilon$ in these high-density regimes. Although these artefacts are reduced when employing a full 3D hand-off, a series of experiments has revealed that reconstructing the $T$ instead of the $\epsilon$ during evolution leads to significant improvements of this issue and is effective across various EOSs.
Finally, for the metric solver, the full-GR Z4c scheme in FIL with the Antelope code typically updates at every sub-step of the time integrator, while in BHAC+, the metric solver is set to update every $0.5~{}M_{\odot}$ for the postmerger simulations (see Appendix C). Both codes use a third-order Runge-Kutta method for time integration and employ a Courant-Friedrichs-Lewy (CFL) factor of $\mathcal{C}_{\rm CFL}=0.2$. Unless otherwise specified, these parameters and numerical methods are the default setups for both codes.
III.2 Hand-off procedure of three-dimensional data
As mentioned above, the hand-off procedure is a key aspect of our approach and allows us to leverage FIL’s accuracy during the inspiral phase and BHAC+’s efficiency, low-memory usage and AMR flexibility in the postmerger phase. The hand-off procedure was first presented in Ref. Ng et al. (2024a), but we have here further refined it by employing a memory-efficient data-initialisation method and by incorporating advanced higher-order interpolation schemes. Both of these improvements are detailed below.
III.2.1 Memory-efficient data initialisation
A non-trivial challenge to face when importing data from FIL into BHAC+ in full-fledged and high-resolution AMR simulations is the different strategies adopted by the two codes in handling the mesh refinements. More specifically, while FIL employs box-in-box refinement, BHAC+ has more flexibility of shaping the refinement boundaries offered by the block-based quadtree-octree AMR. To cope with these differences, we adopt a strategy to reshape the boundaries at each refinement level based on values of the rest-mass density so that the boundary of a refinement level is set by an iso-contour of $\rho$. This solution provides a natural criterion reflecting the distribution of matter and leads to a significant memory saving. Additionally, the finest refinement level is shaped to encompass the dominant rest-mass contribution of the BNS system. With this grid-structure adjustment, each data-point in BHAC+ must be interpolated from the original FIL data.
Another aspect of the hand-off we had to resolve was the reading procedure. In particular, unlike other approaches where an intermediate restart file is generated by a standalone script, e.g., see Lopez Armengol et al. (2022), we directly read and interpolate the data output from FIL within BHAC+, thereby reducing the amount of different software packages employed. This represents a challenge especially in importing the high-resolution data, as additional memory is required to store the new data from FIL and avoid frequent file-read operations. We recall that the data output from FIL consists of multiple HDF5 files containing distinct patches of the simulation domain and including the ghost zones. Hence, when considering a given refinement block in BHAC+, we first determine the number of FIL HDF5 files required to cover such a block in BHAC+. Subsequently, only the relevant FIL-output files are read and the pertinent data are interpolated onto the BHAC+ grid within each refinement block. Besides, in order to save memory, which is essential in high-resolution simulations, we de-allocate the data segments read from FIL after finishing initialisation of one refinement block and continue with loading new pertinent data for the next refinement block. Finally, we should also note that this process is performed in parallel on different processors for different refinement blocks, which makes the hand-off procedure highly efficient. Overall, this method balances the file-reading operations with the memory footprint of the hand-off, and greatly benefits from the parallel nature of this process.
III.2.2 High-order interpolation for spacetime and matter quantities
At hand-off, the rest-mass density $\rho$, proton fraction $Y_{p}$, specific internal energy $\epsilon$, Eulerian velocity $v^{i}$, and conformal factor $\psi$ need to be interpolated from the gridpoint values in FIL to those in BHAC+ that, as discussed above, do not coincide in general. The conformal factor from FIL is used to calculate the gauge-independent conserved quantities, given by $\sqrt{\gamma/\hat{\gamma}}\left(D,S_{j},\tau,DY_{p},B^{j}\right)$. Once Eqs. (2)–(5) are solved, both the metric and the primitive variables are updated Ng et al. (2024a). At the same time, the gauge-independent conserved quantities are kept unchanged during metric initialisation, and we refer readers to Ng et al. (2024a) for a detailed metric interpolation procedure. This approach has been shown to reduce large initial perturbations in the hand-off, which can otherwise cause inconsistencies in the subsequent evolution Ng et al. (2024a). Consequently, any differences in the primitive variables between FIL and BHAC+ arise only from interpolation error and gauge differences. Furthermore, after significant experimentation we have found that the best choice to minimise differences in various global quantities is to employ a third-order Lagrangian interpolation for the metric and the fluid variables.
Special attention is required when importing the magnetic field, for which two approaches are possible. We recall, in fact, that while FIL evolves the vector potential $A^{i}$, BHAC+ solves the induction equation directly in terms of the magnetic field components $B^{i}$. Hence, it is in principle possible to either interpolate directly the values of the magnetic field from FIL as any other fluid variable, or to actually interpolate the vector potential and then compute the magnetic field in BHAC+ by taking the curl operator. The first approach preserves the magnetic-field structure but does not guarantee it to be divergence-free. In contrast, the second method, while introducing small structural differences in the magnetic field due to metric disparities between FIL and BHAC+, guarantees a divergence-free field in BHAC+. We have obviously implemented both approaches and concluded that importing the vector potential provides the most accurate transfer of data. This is because the calculation of the magnetic field from the vector potential yields a divergence-free constraint enforced at machine precision and this is best suited for the constrained-transport approach employed in BHAC+ (Olivares et al., 2019) that benefits from data that is as divergence-free as possible.
At the same time, since a Lagrangian interpolation does not guarantee continuity in the first derivatives and since a curl operator needs to be applied to the vector potential to derive the components of the magnetic field, a different high-order interpolation approach needs to be employed for the vector potential. Also in this case, after extensive experimentation of different approaches and orders, we have concluded that a third-order Hermitian interpolation for the vector potential yields the needed accuracy and smoothness in the first derivatives at a reasonable computational cost.
Tests Baryonic mass EM energy Internal energy
PM-A $0.35\%$ $1.97\%$ $0.31\%$
PM-B $0.29\%$ $0.89\%$ $0.28\%$
PM-C $0.27\%$ $0.85\%$ $0.24\%$
PM-D $0.21\%$ $0.87\%$ $0.20\%$
Table 1: Absolute values of the relative difference of various volume integrals in BHAC+ and FIL at different hand-off stages.
The results of this admittedly elaborate interpolation procedure from FIL to BHAC+ are summarised in Table 1, where we compare the relative differences between FIL and BHAC+ in various global quantities computed via volume integrals, i.e., the baryon mass, the EM energy and the internal energy. Note also that the table contains data relative to different hand-off times as these also play a role in determining the precision of the interpolation. This is because the assumption of conformal flatness becomes increasingly more accurate as the evolution proceeds (see also a discussion in Appendix E) so that the solutions in FIL and BHAC+ are very similar. Notwithstanding the improvement that comes with hand-offs at later times, the largest observed relative difference in baryon mass [see Eq. (17) for the definition] reported in Tab. 1 is $0.35\%$. On the other hand, when considering a point-wise comparison between FIL and BHAC+ quantities we find that the large majority of grid-points exhibit a relative difference of less than $1\%$. Such relative differences can be larger in the presence of strong shock waves, which are frequently formed in the less dense outer regions of the merger remnant, but do not have an impact on the dynamics of the high-density bulk flow. Indeed, the relative differences in the high-density regions are $\lesssim 0.3\%$ for all hand-off times and are mostly due to differences in the field variables between FIL and BHAC+.
HO time $|B^{\rm max}_{0}|$ (FIL) Res. (FIL) Res. (BHAC+) $|h|^{2}$
$[{\rm ms}]$ $[{\rm G}]$ $[{\rm m}]$ $[{\rm m}]$ $[|h|^{2}_{\rm mer}]$
short-term
PM-A 7.57.5~{}7.57.5 $3.7\times 10^{15}$ $300$ $300$ $1/4$
PM-B $13.5$ $3.7\times 10^{15}$ $300$ $300$ $1/8$
PM-C $19.1$ $3.7\times 10^{15}$ $300$ $300$ $1/16$
PM-C-RR $19.1$ $3.7\times 10^{15}$ $300$ $300$ $1/16$
PM-D $30.0$ $3.7\times 10^{15}$ $300$ $300$ $1/32$
PM-D-RR $30.0$ $3.7\times 10^{15}$ $300$ $300$ $1/32$
long-term
LowB-LR $20.0$ $2.0\times 10^{16}$ $300$ $300$ $1/18$
LowB-HR $20.0$ $2.0\times 10^{16}$ $300$ $501$/181181/181 / 18
HighB-LR $20.0$ $1.0\times 10^{17}$ $300$ $300$ $1/18$
HighB-HR $20.0$ $1.0\times 10^{17}$ $300$ $501$/181181/181 / 18
Table 2: List of the simulations performed in this work. In the short-term simulations, “RR” refers to evolutions including the RR corrections. In the long-term simulations, “LR” and “HR” denote the low and high-resolution cases, respectively. HO time is the hand-off time moment. Finally, the last column reports the fraction of the GW amplitude at which the hand-off is performed.
III.3 Initial data and simulation setups
The initial data is computed using the TNTYST Togashi et al. (2017) EOS and refers to equal-mass and irrotational binaries with a total ADM mass of $2.55~{}M_{\odot}$, initialised at a separation of $45\,\mathrm{km}$. When the separation between the “barycentres” of the two stars is $\lesssim 13.3\,\mathrm{km}$, a purely poloidal magnetic field Etienne et al. (2015); Chabanov et al. (2023) is seeded inside each neutron star. The seed magnetic field is given by
$A_{i}=A_{\rm b}\left[-(x^{j}-x^{j}_{\rm NS})\epsilon_{ij}\right]\,\max\{p-p_{\rm cut},0\}\,,$ (15)
for $i=x,y$ and $A_{z}=0$, where $A_{\rm b}$ is a constant chosen to adjust the magnetic-field strength, $\epsilon_{ij}$ is the Levi-Civita symbol, $x_{\rm NS}$ denotes the coordinate center of the neutron star, $p$ is the fluid pressure, and the cut-off pressure is set to $p_{\rm cut}=1.0\times 10^{-8}$. Three different strengths for the seeding magnetic field are used in this work: $3.7\times 10^{15}\,\rm G$, $2.0\times 10^{16}\,\rm G$, and $1.0\times 10^{17}\,\rm G$. The corresponding EM energies at the seeding time are $7.6\times 10^{47}$, $2.0\times 10^{49}$, and $5.4\times 10^{50}\,\rm erg$, respectively. A summary of the simulations performed in this work together with essential information on the hand-off is presented in Tab. 2, where “RR” in the simulation name denotes the simulation with RR scheme, while “LR” and “HR” refer to the low and high resolution cases, respectively.
Both codes adopt 3D Cartesian coordinates with $z$-symmetry and employ a refinement ratio of $2$ with $6$ refinement levels, where the finest level has a boundary at $24\,\mathrm{km}$ and a resolution of $300\,\mathrm{m}$. For the simulations reported here, FIL has a computational grid with outer boundaries at $1500\,\mathrm{km}$ in all three spatial directions. In contrast, BHAC+ adjusts the grid setup by shrinking the outer boundary to $600\,\mathrm{km}$, focusing on the properties of the postmerger remnant. Note that unlike a full-GR scheme, the xCFC scheme does not require a large domain to apply outgoing radiative boundary conditions or to ensure sufficient damping of the constraint violation. Additionally, each refinement level in BHAC+ has a spherical structure that more accurately conforms to the shape of the merger remnant than a Cartesian box used in FIL.
Figure 1: Evolution of the GW amplitude. Indicated with different dashed lines are either the time of merger or the different times when the hand-off is performed.
IV Validation of the hand-off and RR
The CFC approach has been successfully applied in the study of uniformly and differentially rotating neutron stars Cook et al. (1996); Iosif and Stergioulas (2014, 2021); Cheong et al. (2024, 2025), with only minor deviations observed compared to the full-GR solutions. Furthermore, our previous work Ng et al. (2024a) has demonstrated that the CFC approach can closely approximate the full-GR solution in the highly nonlinear system of neutron stars head-on collisions in full 3D simulations. However, the performance of the CFC in realistic BNS postmerger scenarios with magnetic fields and its comparison with full-GR simulations, has not been explored thus far.
In this Section, we compare the evolution after hand-off in BHAC+ with the corresponding evolution of the postmerger remnant in FIL for four different hand-off times. In particular, after defining the time relative to the merger as $\bar{t}:=t-t_{\rm mer}$, we denote the time of hand-off as $\bar{t}_{{}_{\rm HO}}$ and consider the transfer of data from FIL to BHAC+ at four different values of $\bar{t}_{{}_{\rm HO}}$, i.e., from PM-A to PM-D (this information is also summarised in Tab. 2 and can be appreciated in Fig. 1). Furthermore, to make the hand-off times as robust and reproducible as possible, we set them on the basis of the strength of the GW radiation emitted by the postmerger remnant. More specifically, we hand-off the simulation from FIL to BHAC+ when the amplitude of the emitted GWs decreases to $1/4$ (PM-A), $1/8$ (PM-B), $1/16$ (PM-C), and $1/32$ (PM-D) with respect to its peak value (see also see Tab. 1). Hereafter, special attention will be paid on the results of the hand-off made at the “late” time of $\bar{t}=30\,\rm ms$, i.e., PM-D.
IV.1 On the role of hand-off time
In Fig. 2, we report the evolution of three representative fluid quantities as simulated by BHAC+ with varying $\bar{t}_{{}_{\rm HO}}$ values and compare them with the corresponding results of FIL, both when the RR terms are included or ignored. Starting from the top panel of Fig. 2, which reports the maximum gauge-invariant conserved density $D^{*}_{\rm max}$, it is possible to note that high-frequency oscillations are introduced immediately following hand-off, though they are quickly smoothed out by subsequent evolution within $5~{}\mathrm{ms}$. The perturbation amplitudes remain comparable across all BHAC+ cases, suggesting that they primarily originate from mapping errors rather than from gauge differences.
Figure 2: Evolution of the maximum value of the conserved rest-mass density $D^{*}$ (upper panel), of the total EM energy (middle panel), and of the toroidal/poloidal EM energies (bottom panel). Different lines refer to the values computed by FIL (red solid line) and those by BHAC+ at different hand-off time cases, with and without RR (blue-shaded solid lines). Note that the evolutions PM-C and PM-D nearly overlap with the corresponding evolutions that include RR terms.
The evolution in FIL reports a well-known dynamics, namely, that the merger remnant reaches a quasi-equilibrium state after emitting energy and transferring angular momentum, primarily through GW emission and ejecta, and ultimately settling as a differentially rotating massive star with a gradually increasing central density (see Kastaun and Galeazzi (2015); Hanauske et al. (2017); Iosif and Stergioulas (2022); Cassing and Rezzolla (2023) for a discussion of the differential rotation profiles). At the earliest hand-off time (PM-A with $\bar{t}_{{}_{\rm HO}}=7.5~{}\mathrm{ms}$), a rapid contraction occurs in the remnant because the GW amplitude only decrease to $1/4$ of its peak value. In this case, the use of xCFC without RR introduces a notable relative difference of approximately $4.9\%$ in $D^{*}_{\rm max}$ as it reaches a quasi-equilibrium state at $\bar{t}=25.2~{}\mathrm{ms}$. However, as the GW emission from the merger remnant decreases, at $\bar{t}=40\,\mathrm{ms}$, the results from BHAC+ align well with FIL even in cases without RR, with relative differences of $2.1\%$, $1.6\%$, and $1.4\%$ in $D^{*}_{\rm max}$ for the cases PM-B, PM-C, and PM-D, respectively.
Although all evolutions exhibit a gradual increase in $D^{*}_{\rm max}$, those cases that do not include RR generally show slightly lower values simply because the energy of the remnant does not have the possibility of leaving the system and hence of reaching higher compactnesses, as noted in Ng et al. (2024a). At the same time, the inclusion of RR terms, as for cases PM-C-RR and PM-D-RR, does improve the expected behaviour and leads to evolutions with a gradual increase in the maximum rest-mass density starting from right after the hand-off and ultimately reaching a quasi-equilibrium state similar to FIL, with relative differences of $0.32\%$ and $0.74\%$ in $D^{*}_{\rm max}$ at $\bar{t}=40~{}\mathrm{ms}$, respectively. Overall, these simulations demonstrate that the inclusion of RR terms is important to reach a postmerger evolution that is as faithful as possible.
The middle panel of Fig. 2 reports the evolution of the EM energy [see definition in Eq. (19)] and illustrates that the magnetic-winding process is very well captured by BHAC+ in cases when the hand-off is done when the GW amplitude decreases below $1/16$ of its peak value, i.e., cases PM-C and PM-D. This accurately represents both the rate of change in EM energy and the time at which winding ceases. Indeed, the largest relative differences in total EM energy is of only $4.7\%$ for PM-D, but obviously increases to $15.4\%$ and $31.0\%$ when the hand-off is done earlier, i.e., for cases PM-C and PM-B, respectively.
Finally, in the bottom panel of Fig. 2 we report the evolution of the toroidal and poloidal EM energies [see definition in Eq. (20)]. As well-known, the toroidal magnetic field dominates the postmerger magnetic energy as a result of the KHI and the winding process, which mostly convert kinetic energy into toroidal magnetic energy. On the other hand, the poloidal EM energy continues to dissipate until the winding process ends, due to a combination of shocks and violent oscillations within the remnant, with a mild increase at later times. Interestingly, oscillations in the evolution of poloidal EM energy emerge in BHAC+ if the hand-off is performed too early. These oscillations reflect the development of an $\ell=2,m=1$ barmode instability East et al. (2015); Lehner et al. (2016); Radice et al. (2016); East et al. (2019); Papenfort et al. (2022); Topolski et al. (2024) triggered by the error introduced at hand-off if this is made too early. Indeed, the amplitude of the oscillations is reduced for later hand-offs (the smallest amplitude is observed in PM-D, due to greater axisymmetry and conformal flatness (see also a discussion in Appendix E) at this stage) and it decays rapidly with time in all cases. More importantly, because these oscillations only affect the poloidal EM energy and the latter is about two orders of magnitude smaller than the toroidal one, the overall evolution is hardly affected (see middle panel).
In summary, our results show that using the GW amplitude represents a gauge-invariant, model-independent, and essentially resolution-independent manner to assess when the hand-off can be performed. In addition, we suggest that performing the hand-off when the GW amplitude has reached $\simeq 1/16$ of the peak amplitude represents a robust choice. While these conclusions are based on a limited number of tests, the underlying motivation is clear: the smaller the GW emission, the better the CFC approximation and its use at hand-off. At the same time, we also recommend that some experimentation is made around the hand-off time. While this experimentation brings in some extra computational costs, the latter are minute when compared with those associated with the long-term evolution that can be performed under optimal hand-off conditions.
IV.2 Case PM-D: xCFC vs full-GR results
As discussed above, the hand-off at $\bar{t}_{{}_{\rm HO}}=30\,\rm ms$ provides the best match with the full-GR evolution computed by FIL, thus making it worth a closer investigation. Hence, we start our comparison between the results of BHAC+ (solid lines) and FIL (dashed lines) by presenting in Fig. 3 the azimuthally averaged profiles of the rest-mass density and of the angular velocity on the equatorial plane of the merger remnant at different times after hand-off. The differential-rotation profiled observed (top part of Fig. 3) is well known Kastaun and Galeazzi (2015); Hanauske et al. (2017); Cassing and Rezzolla (2023) and shows a peak at $r\approx 8\,\rm{km}$, a Keplerian fall-off of the type $r^{-3/2}$, and gradually flattens over time due to winding and dissipation in both codes Duez et al. (2004, 2006); Ciolfi et al. (2019) (see also the discussion in Sec. V.2). A measure of the very good match between the two evolutions can be found in the relative difference in the rotation frequency, which remains below $2\%$ ($5\%$) at $20\,\rm ms$ ($30\,\rm ms$) after the hand-off. Similarly, the difference in the location of the angular-velocity maximum between the two codes remains within $1\%$ across all the times considered, indicating that angular momentum redistribution and frame-dragging effects are similarly captured in both codes at this stage.
Figure 3: Azimuthally averaged profiles of the angular velocity (upper panel) and of the rest-mass density (lower panel) as computed at different times by FIL (dashed lines) and by BHAC+ (solid lines). The data is computed on the equatorial plane and refers to the case PM-D, i.e., with a hand-off at $\bar{t}_{{}_{\rm HO}}=30\,\rm ms$.
The different profiles of the rest-mass density produced by the two codes (bottom part of Fig. 3) are essentially identical in the core regions within $r\approx 8\,\rm{km}$ and do not vary significantly at least within the first $30\,{\rm ms}$ following the hand-off. In the outer regions of the remnant, matter expands outward through oscillations, leading to an increase in rest-mass density over time for both codes. The largest relative difference remains below $1.5\%$ within $r\approx 8\,\rm km$ (where $96.4\%$ of the rest-mass is concentrated) and still remains below $10\%$ within $r\approx 15\,\rm km$ (where $97.2\%$ of the rest-mass is concentrated). Larger differences appear at large radii but these are related to the slightly different locations of the spiral shocks and involve amounts of matter that are not dynamically relevant.
Figure 4: Projections on the three principal planes of the temperature and Lorentz factor (left panel) and of the toroidal and poloidal magnetic-field strengths (right panel) for the merger remnant from case PM-D at $20\,\rm ms$ after hand-off. For each plane, the left portion refers to simulations from BHAC+ and the right one from FIL. Note that the negative-$z$ regions are included for visualization purposes only, as the simulations assume symmetry across the $z=0$ plane.
[…]
Similar considerations can be made also for the distributions of the poloidal and toroidal magnetic field (right panel of Fig. 4). In this case, the most significant feature is a strong toroidal magnetic field forming a ring structure located between $\sim 3.4\,\rm km$ and $\sim 6.2\,\rm km$ due to winding caused by strong differential rotation. This prominent structure is reproduced in both codes and the match between the two is very good, despite the use of different methods for evolving the magnetic field Etienne et al. (2015); Most et al. (2019); Olivares et al. (2019). In the low-density regions (i.e., $\rho\ll\rho_{\rm sat}$), the magnetic field is dominated by turbulence and we do not expect the distributions of magnetic field in the two codes to have a significant resemblance. Yet, the statistical properties of the turbulent magnetic field in the two codes is very similar, as can be easily appreciated in the right panel of Fig. 4. We also note that, as mentioned above, the toroidal magnetic-field strength is systematically stronger than that of the poloidal magnetic field on all three principal planes (see also the bottom panel of Fig. 2). For example, on the equatorial plane, the maximum values are $2.6\times 10^{16}\,\rm G$ and $3.2\times 10^{15}\,\rm G$ for the toroidal and poloidal components, respectively. As discussed in several works in the literature Ciolfi et al. (2019); Palenzuela et al. (2022); Chabanov et al. (2023); Kiuchi (2024); Musolino et al. (2024a), this difference between the poloidal and toroidal magnetic fields arises for two reasons: first, the winding process significantly amplifies the toroidal magnetic field; second, the lack of an efficient dynamo process that limits the conversion of toroidal to poloidal magnetic field in the current simulations.
In summary, the results discussed in this Section provide very strong evidence that BHAC+, despite a simpler description of the gravity sector, namely, the use of the CFC approximation reproduces the full-GR results coming from FIL extremely well in all fluid quantities. In addition, the inclusion of RR terms in the xCFC scheme further improves the accuracy. Furthermore, while these considerations have been deduced when considering the “late” hand-off at $\bar{t}_{{}_{\rm HO}}=30\,\rm ms$, PM-D (when the GW amplitude has reached $\sim 3\%$ of its peak value), similar considerations can be made also with earlier hand-offs (see Appendix D for a similar discussion in the case of the hand-off PM-C).
As a final remark, we note that the differences reported here in the postmerger evolutions from FIL and BHAC+ are significantly smaller than those reported in Refs. Espino et al. (2023); Neuweiler et al. (2024) in terms of the evolution of the maximum rest-mass density, total EM energy, magnetic-field strength, rotational profiles, and remnant lifetimes among different GRMHD codes. Despite all of these codes employed in Refs. Espino et al. (2023); Neuweiler et al. (2024) make use of full-GR treatments and adopt similar numerical schemes, differences remain and underline the difficulties in providing very accurate descriptions of the postmerger evolution.
V Impact of magnetic-field strength on long-lived “magnetars”
While the previous Sections have been dedicated to the discussion and testing of the strategy developed to optimise the hybrid strategy proposed in Ref. Ng et al. (2024a) to perform long-term simulations of postmerger remnants, the remainder of the paper is dedicated to use the FILBHAC+ infrastructure to actually explore a specific issue on the evolution of the postmerger remnant, namely, the impact of the magnetic-field strength on the dynamics and EM emission from a long-lived highly magnetized merger remnant, or “magnetar”, produced in a BNS merger.
We recall that the magnetic-field strengths in the BNS merger remnant are expected to reach up to equipartition values of $\sim 10^{16}$-$10^{17}\,\rm G$, given the large energy reservoir in the remnant, both in terms of kinetic and binding energy, and that current resolutions of numerical simulations still can only partially resolve two key amplifying effects, i.e., KHI and MRI, of the magnetic field. This is shown by sub-grid modelling simulations  Aguilera-Miret et al. (2022); Palenzuela et al. (2022), or simulations adding phenomenological dynamo terms Shibata et al. (2021); Most and Quataert (2023); Most (2023), and also very high-resolution simulations Kiuchi et al. (2015a); Chabanov et al. (2023); Kiuchi et al. (2024); Musolino et al. (2024a). Given these limitations and the enormous costs of carrying out simulations with very high resolution and on very long timescales, we here consider a different approach. Namely, we consider and contrast two different scenarios that differ only in the strength of the initial magnetic field, i.e., $2.0\times 10^{16}\,\rm G$ and $1.0\times 10^{17}\,\rm G$, so as to assess what is the role of the amplification mechanism – whatever that is – on the subsequent long-term evolution of the postmerger. Hereafter, we will refer to these two initial magnetic fields as low-magnetic-field (LMF) and high-magnetic-field (HMF), respectively.
More specifically, for each initial magnetic-field strength, we have performed a simulation with FIL with the finest resolution of $300\,\rm m$ of the inspiral, merger and postmerger. Such a solution was then transferred to BHAC+ with a hand-off at $20\,\rm ms$ (this is when the GW amplitude has dropped below $1/16$ of its peak value) and then evolved with two different resolutions of $300\,\rm m$ and $150\,\rm m$ so as to have a postmerger evolution over a timescale of $200\,{\rm ms}$ with a resolution that would be prohibitive with FIL. More specifically, in high-resolution simulations, we add the finest refinement level in the region [16, $[-16,\,16]\,\rm km$, which adequately covers the region where $\rho>10^{12}\,\rm g\,cm^{-3}$. Hereafter, we will refer to these four scenarios respectively as LowB-LR, LowB-HR, HighB-LR, and HighB-HR (see also Tab. 2).
V.1 Impact on the EM-energy evolution
Figure 5: EM-energy evolution for simulations with LMF and HMF. Different line-styles represent different classes of EM energy, namely, $E_{\rm EM,tot}$ (thick solid line), $E_{\rm EM,tor}$ (thick dashed line), and $E_{\rm EM,pol}$ (thin solid line), while different colours indicate simulations that differ either in resolution or initial magnetic field (see Tab. 2 for the nomenclature of the various cases). The black vertical dash-dotted line marks the hand-off time, with the results from BHAC+ (FIL) shown on the right (left) with respect to it. The vertical black (red) dashed line indicates the time at which the remnant undergoes a slow down at the peak of EM energy in the HMF (LMF) scenario.
We start our discussion by reporting in Fig. 5 the evolution of the EM energy in the four merger remnants. Lines of different colours refer to the four scenarios considered and the black vertical dash-dotted line represents the hand-off time, with the results from BHAC+ (FIL) shown on the right (left) with respect to it. Note that, as remarked above, the LowB-LR and HighB-LR simulations in FIL are then also evolved at higher resolution in BHAC+ as LowB-HR and HighB-HR; this explains why there are only two sets of lines to the left of the hand-off time.
Both the toroidal and poloidal EM energy grow rapidly during the KHI stage, peaking at around $\bar{t}\approx 2\,\rm ms$. After this point, the EM energy begins to rapidly decrease, mostly due to the lack of sufficient turbulent vortices to counteract dissipation (see discussion in Chabanov et al. (2023)). As the merger remnant develops a large-scale differential rotation structure, the winding stage commences and the toroidal magnetic-field strength exhibits the characteristic linear growth over time, which can last up to $\bar{t}\simeq 30\,{\rm ms}$ ($\bar{t}\simeq 90\,{\rm ms}$) for the HMF (LMF) case at high resolution. Over the same period of time, we observe a rapid decay in the poloidal-energy component, primarily due to numerical dissipation caused by a combination of shocks and violent oscillations within the remnant.
Magnetic winding ceases when the toroidal magnetic fields become strong enough to exert an effective torque on the fluid, resulting in angular momentum redistribution. This transition, which we denominate as “slow-down” occurs earlier in the HMF case (i.e., $\bar{t}\simeq 30\,{\rm ms}$) than in the LMF case ($\bar{t}\simeq 90\,{\rm ms}$); the relevant times are marked with vertical dashed lines in Fig. 5. It is important to underline that the slow-down does not mark the end of the magnetic-field winding, which indeed continues also afterwards, but rather the beginning of a different regime in which winding of the poloidal magnetic field into the toroidal one still takes place but is slowed down, i.e., it leads to a sublinear in time growth, because of the intense back-reaction of the magnetic tension (see also the discussion in Sect. V.2). In addition, we should note that we expect the specific threshold depends, in addition to the initial magnetic-field strength, also on the rotational profile of the merger remnants and hence also on the mass ratio and EOS. This redistribution of angular momentum simultaneously generates an effective poloidal fluid motion, counteracting the dissipation of poloidal EM energy. Additionally, there is another contribution to the poloidal EM-energy growth after the EM amplification ceases and it comes from buoyant toroidal field lines that rise to higher latitudes, causing magnetic flares that build structured poloidal magnetic fields in the polar region (see discussion below).
An important and interesting behaviour found is that when the initial magnetic field is not large, i.e., in the LMF case, the winding period lasts longer and achieves higher EM energy than the HMF case. Specifically, the high-resolution simulations show peak EM energies of $2.7\times 10^{51}\,\rm erg$ for the HMF case and $7.1\times 10^{51}\,\rm erg$ for the LMF case, hence with an EM energy that is about three times larger. Naively, one would expect that starting with a weaker magnetic field would also yield a lower final magnetic field, which is the opposite of what our simulations have revealed. The explanation of this counter-intuitive behaviour has to do with the nonlinearity of the amplification process and on the ability to use the available kinetic energy stored in the fluid to convert it into magnetic energy. In particular, these results show that when the initial magnetic strength is large, it exerts a substantial feedback on fluid motion, suppressing small-scale turbulence through magnetic tension, thus preventing winding and effectively weakening the amplification very rapidly or, equivalently, leading to an early slow-down (the simple linear winding essentially stops around $\bar{t}\approx 30\,\rm ms$ for the HMF case). On the contrary, when the initial magnetic strength is small, the fluid is less constrained in its dynamics and can be used for a larger number of winding periods, thus leading to a stronger final magnetic field and a delayed weakening of the amplification (the linear winding essentially stops around $\bar{t}\approx 90\,\rm ms$ for the LMF case). This interpretation of the dynamics is also supported by the analysis of the fraction of the kinetic energy that has been converted into EM energy, which is only $\simeq 2\%$ in the HMF case and $\simeq 6\%$ in the LMF case. Finally, we note that the use of higher resolution has the effect of prolonging the winding period, delaying the appearance of the slow-down time; this difference is of $\sim 6\,\rm ms$ or $20\,\rm ms$ in the HMF case and LMF cases, respectively. This extension of the winding can be explained by the fact that higher resolution naturally reduces numerical resistivity, thus allowing for more efficient bending of the magnetic field lines via differential rotation and hence to a more efficient winding (see also the discussion in Refs. Dionysopoulou et al. (2015); Shibata et al. (2021)).
Figure 6: Spacetime diagram of azimuthally averaged sections of the angular-velocity profile on the equatorial plane of the remnant for the LMF (left panel) and the HMF scenarios (right panel). The transparent region reports the result of FIL, while the rest of the diagram reports BHAC+ data. The horizontal dashed white lines represent the time of slow-down, while the dashed, dot-dashed, and solid black lines report the worldlines of rest-mass density contours at $10^{13},10^{13.5},\text{and}\,10^{14}\,\rm g~{}cm^{-3}$, respectively.
V.2 Impact on the differential-rotation evolution
It is well-known that the interaction between differential rotation and magnetic fields in stars inevitably leads to the generation of magnetic field, but also to the suppression of differential rotation (see Refs. Rezzolla et al. (2000); Shapiro (2000); Rezzolla et al. (2001); Duez et al. (2006), and also Refs. Duez et al. (2004); Radice (2017) for an equivalent process in the presence of shear viscosity). It is therefore interesting to investigate how the process of removal of differential rotation takes place in a BNS remnant and how this depends on the (initial) strength of the magnetic field.
To this scope, we report in Fig. 6 a spacetime diagram over the whole timescale of the simulation of azimuthally averaged sections of the angular-velocity profile on the equatorial plane of the remnant for the LMF (left panel) and the HMF scenarios (right panel), respectively. Reported in the white-shaded area are the results from FIL, while the data after hand-off refers to the evolution obtained with BHAC+. In such a spacetime diagram it is particularly interesting to contrast the early part of the evolution at $\bar{t}\sim 25\,{\rm ms}$ with the final ones at $\bar{t}\sim 200\,{\rm ms}$. Clearly, the remnant has a significant amount of differential rotation at first (see also Fig. 3) but this is gradually changed by the interaction between the rotation of the fluid and the magnetic-field tension. As a result, in both scenarios of low and high initial magnetic field, the degree of differential rotation is significantly smaller and differential rotation is essentially removed in the HMF case by the time the simulation is ended.
Figure 7: Spacetime diagrams of the Parker-instability criterion (left panel) and of the inverse plasma-$\beta$ (right panel) averaged on the $(x,z)$ plane in simulations with initial LMF (left portions) and HMF (right portions), respectively. The top part of the figure refers to the “funnel” region, i.e., the region with a polar angle less than $\pi/6$, while the bottom part refers to the “disk”, which is the remainder. The breakout times are shown with horizontal white solid lines for each simulation. Vertical dashed, dash-dotted, and solid cyan lines represent the worldlines of rest-mass density contours at $10^{13},10^{13.5},\text{and}\,10^{14}\,\rm g~{}cm^{-3}$, respectively.
Note also the strong correlation that we find between the onset of the slow-down (i.e., the quenching of the linear-in-time winding discussed above) and the transition to process of removal of differential rotation. This is not surprising, since the slow-down marks the time when linear-in-time winding ceases because magnetic tension is exerting a sizeable torque on the fluid and opposing the generation of additional toroidal magnetic field. Another interesting aspect of the suppression of differential rotation that can be appreciated from the spacetime diagram in Fig. 6 is that the remnant also undergoes a significant redistribution of angular momentum, which can either increase the angular velocity in the core (LMF case) or decrease it. These structural changes in the angular momentum also show that the initial peak in differential rotation is not simply washed out. Rather, before the angular velocity attains an almost uniform rotation profile (see the HMF case in Fig. 6 at $\bar{t}\gtrsim 150\,{\rm ms}$) two different peaks appear at smaller and larger radii, which are progressively erased (this can be seen also in the LMF case in Fig. 6 but takes place much later). Hence, the “forking” of the angular-velocity maximum can also be taken as a good proxy of the loss of efficiency in the winding and the beginning of the slow-down phase. In addition, the redistribution of angular momentum can also be tracked via the worldlines of the rest-mass density at $10^{13}$ (black dashed line), $10^{13.5}$ (black dot-dashed line), and $10^{14}\,\rm g~{}cm^{-3}$, which indicate that all of the corresponding iso-contour are moving inwards as a result of the depletion of angular momentum in the remnant and that is transported outwards as a result of the MRI and winding.
Overall, our results are in good agreement with what has been observed in simulations with added effective shear Fujibayashi et al. (2020), or resistivity Dionysopoulou et al. (2015); Shibata et al. (2021), and with very strong magnetic fields Bamber et al. (2025). When taken together, this bulk of works suggests that differential rotation in a BNS remnant is likely to be washed out over a timescale that clearly depends on the strength of the magnetic field but that can be reasonable narrowed between $\bar{t}\sim 200\,{\rm ms}$ for magnetic-fields of the order of $\simeq 10^{17}\,{\rm G}$ and $\bar{t}\sim 300\,{\rm ms}$ for weaker magnetic-fields of $\simeq 10^{16}\,{\rm G}$. However, we caution that when starting from realistic initial-field strengths of $\simeq 10^{11}\,{\rm G}$, the winding process might be prolonged due to the turbulent magnetic field in the early postmerger phase.
V.3 Parker instability and flares
As mentioned in Ref. Musolino et al. (2024a), the physical conditions in the outer layers of the merger remnant and close to the polar axis can lead to a global breakout of the plasma as a result of the development of the Parker instability. We recall that the latter is described by the criterion Parker (1966)
$\mathcal{P}:=\frac{d\log{p}}{d\log{\rho}}-1-\frac{\beta^{-1}\left(1+2\beta^{-1}\right)}{2+3\beta^{-1}}\,,$ (16)
where $\beta:=p/(b^{2}/2)$ is the ratio of the fluid pressure to the magnetic pressure. If the magnetic-field pressure continues to increase within a region, the third term in Eq. (16) becomes dominant and once the condition for instability $\mathcal{P}<0$ is met, low-density matter with high total pressure will rise due to buoyancy, carrying the magnetic field along with it. We refer to this process as to the “breakout” as it characterises a sharp transition between two different states of the plasma in the polar region.
Figure 8: Magnified section on the $(x,z)$ plane of the merger remnant at $\bar{t}=60\,\rm ms$ (top row) and $\bar{t}=170\,\rm ms$ (bottom row) so as to highlight the generation of the magnetic eruptions (flares). Starting from the left, the different columns report: the Parker instability criterion $\mathcal{P}$, the inverse plasma-$\beta$, the specific entropy $s$, and the Lorentz factor $W$. Magnetic-field lines and fluidlines are shown in the first and third columns, respectively. For each subpanel, the dotted, dash-dotted, dashed, and solid cyan lines show the rest-mass density contour at $10^{11},10^{12},10^{13},\text{and}\,10^{14}\,\rm g~{}cm^{-3}$, respectively.
To analyze the occurrence of the Parker instability, we use the approach proposed in Ref. Musolino et al. (2024a) to divide the simulation domain into two parts: the region with a polar angle less than $\pi/6$, defined as the “funnel”, and the remainder, defined as the “disk”. Figure 7, in particular, reports the spacetime diagrams of the Parker-instability criterion (left panels) and of the inverse plasma-$\beta$ (right panels) for simulations with different magnetic-field strengths. Both quantities are averaged over the polar angle on the $(x,z)$ plane. While the top plots refer to the funnel region, the bottom ones show the disk region; at the same time, for each plot, the left and right portion refers to the LMF and HMF case, respectively. The breakout time, i.e., the time when $\mathcal{P}$ changes sign, is shown with horizontal white solid lines for each simulation, while vertical dashed, dash-dotted, and solid cyan lines represent the worldlines of rest-mass density contours of $10^{13},10^{13.5},\text{and}\,10^{14}\,\rm g~{}cm^{-3}$, respectively.
Using the spacetime diagrams, it is then straightforward to recognise that a breakout clearly takes place in the funnel region (top panels in Fig. 7) and that this is true for the LMF and the HMF scenarios. What varies between these two cases is mostly the time of the breakout, which obviously takes place earlier for the HMF scenario ($\simeq 25\,\rm ms$) than in the LMF one ($\simeq 45\,\rm ms$); the breakout also shows signs of intermittency in the case of low magnetic fields, e.g., at $\bar{t}\simeq 110$ or $190\,\rm ms$, mostly because in this case the stability criterion (16) is only mildly violated. Also different is the strength of the breakout transition, which is clearly more marked in the case of stronger initial magnetic fields (right portions in Fig. 7). This behaviour has a clear physical interpretation: given a rather similar dynamics of the matter, stronger magnetic fields will produce earlier the conditions for the buoyancy of the low-density matter from the surface of the HMNS. What discussed so far applies also to the disk region (bottom panels in Fig. 7), although in these scenarios the breakout is only marginal in the case of the LMF case and is comparatively weaker in the case of the HMF scenario. A phenomenology of this type can be easily interpreted when considering that the fluid pressure in the disk is considerably larger than in the funnel and hence it is harder for the fluid to breakout from the disk region of the HMNS. We will further discuss the impact of this breakout in terms of the observable EM emission in Sec. V.4.
After breakout, the outer layers of the HMNS, i.e., with $\rho\lesssim 10^{12}\,\rm{g~{}cm^{-3}}$, also experience local eruptions of plasma, or “flares”. These can be appreciated both from the spacetime diagrams in Fig. 7 and in Fig. 8, which report vertical sections on the $(x,z)$ plane of the Parker criterion, the inverse plasma-$\beta$, the fluid specific entropy [^2] We here use the specific entropy and not the temperature as it provides a better proxy of the heating in the low-density plasma., and the Lorentz factor; for each panel, the left and right portions refer to the LMF and HMF scenarios, respectively.
Figure 9: The same as in Fig. 8, but showing the large-scale structure at time $\bar{t}=170\,\rm ms$.
These flares, which have been reported also in other works and scenarios Carrasco et al. (2019); Nathanail et al. (2020, 2021b); Most and Philippov (2020); Most and Quataert (2023); Mahlmann et al. (2023); Musolino et al. (2024a), appear not only in the polar regions but also near the equatorial plane, enhancing both mass outflow and angular momentum transport by moving to larger radii matter with large specific angular momentum. Flares appear in both HMF and LMF cases but exhibit distinct characteristics: in the HMF case, flares are stronger and emerge from both low-latitude and polar regions, while in the LMF case, flares are weaker and predominantly emerge from the polar region (see Sec. V.5 for details). These flares help establish structured magnetic fields and clear the funnel region, facilitating the development of highly collimated, quasi-steady outflows (see Sec. V.4 for more on their properties). References Zappa et al. (2023); Musolino et al. (2024a); Ng et al. (2024) suggest that neutrino fluxes are significantly enhanced along the polar directions, partly because the equatorial direction is blocked by the presence of dense matter, and partly because of re-radiation from the disk after neutrino re-absorption. While neutrinos are not essential for the breakout and the subsequent flares Musolino et al. (2024a), their inclusion effectively reduces baryon-loading in the polar regions, leading to more structured, continuous, and strongly collimated outflows. Although the Poynting fluxes in our simulations are less steady than those reported in Ref. Musolino et al. (2024a), and this is mostly due to the absence of neutrinos sweeping the polar funnel, strong Poynting fluxes are still produced in both the LMF and HMF cases (see the discussion in Sec. V.4 below)
Different initial magnetic-field strengths also lead to clear qualitative differences in the flares, as can be appreciated when comparing the left (LMF) and right (HMF) portion of each panel in Fig. 8. More specifically, in the HMF scenario, we first observe that flares are initiated from density regions of $\rho\approx 10^{12}\,\rm g~{}cm^{-3}$ (see the highly distorted magnetic flux-tube that is a sign of violent magnetic buoyancy) and further amplified in the density region of $\rho\approx 10^{11}\,\rm g~{}cm^{-3}$. Besides, they appear both in the disk and in the funnel, where matter can be heated to specific entropies of $\sim 30\,{\rm k_{b}}/\text{baryon}$, much comparatively colder matter is present at lower latitudes. Lastly, we observe some of these low-latitude flares float upward, enhancing the Poynting flux in the polar region, while a smaller fraction propagates outward, contributing to low-latitude emission. By contrast, in the cases with LMFs, flares essentially do not appear in the disk region and are also much weaker in the funnel (compare the left and middle panels in Fig. 8). Besides, the weak flares in the funnel region are originated from density regions far below $\rho\approx 10^{11}\,\rm g~{}cm^{-3}$. These hold true even when the magnetic-field strength within the merger remnant of the LMF case exceeds that of the HMF case, as the strong magnetic field in the LMF case is mostly confined to regions with densities above $10^{14}\,\mathrm{g~{}cm^{-3}}$ before the slow-down stage. It is worth remarking that the comparison that can be made in Fig. 8, and hence the ability to assess the different dynamics that characterize the merger remnant under different magnetic-field strengths, is possible only with a comparative study of the type presented here. Indeed, had we adopted a single value of the initial magnetic field, the picture we would have deduced would have inevitably been a partial one.
By the end of all simulations, and both in the LMF and HMF scenarios, the merger remnant becomes approximately spherical (see density contours in Fig. 8), partly due to the angular-momentum loss and partly due to the reduced pinch-effect of the strong toroidal magnetic field that produces a prolate distribution of matter Frieben and Rezzolla (2012). The reduced deformation of the magnetar is also responsible for a reduction in the amplitude of the remnant oscillations and for the dynamical ejection of matter, both in the polar and equatorial directions. As a result, matter will tend to accumulate, and in the absence of a neutrino wind sweeping away to large distances, it will lead to a change in the stratification in the outer layers of the HMNS and hence to a reduced violation of the Parker instability criterion. In turn, this results into a partial weakening of the Poynting flux around $\sim 200\,\rm ms$ (see Figs. 7 and 11, and the discussion below).
V.4 Properties of the outflow
Figure 10: Distributions of the Poynting fluxes as projected on two-spheres at a coordinate radius of $300\,\rm km$ at $\bar{t}_{1}=100\,\rm ms$ (top row) and $\bar{t}_{2}=170\,\rm ms$ (bottom row), for the low (left panels) and high (right panels) initial magnetic fields (LMF and HMF, respectively). Note that in the case of LMF scenarios, the Poynting luminosity is multiplied by 10 so that it is comparable with the values of the HMF case.
Obviously, the Poynting flux represents one of the best tools to obtain EM information and hence construct a multi-messenger description of the BNS merger system. As a result, they have been explored, in different forms and approximations, for almost a decade (see, e.g., Refs. Liu et al. (2008); Anderson et al. (2008); Rezzolla et al. (2011); Kiuchi et al. (2015b) for some of the initial works). We here discuss how to use the Poynting flux to distinguish between the long-term evolutions with low and high initial magnetic fields. We start by reporting in Fig. 9 the two-dimensional properties of the outflow on the $(x,z)$-plane in terms of the same quantities used in Fig. 8, namely, from left to right, the Parker criterion, the inverse plasma-$\beta$, the specific entropy, and the Lorentz factor. To emphasise the large-scale structure of the collimated outflows, we select $\bar{t}=170\,\text{ms}$, the point at which the Poynting flux in the LMF case reaches its peak across the entire simulation.
Figure 9 clearly shows that highly collimated outflows form in the polar regions of both the LMF and HMF cases and are characterized by ordered and large-scale poloidal magnetic-field lines that serve as guide for the motion of the fluid. These outflows are also characterized by high values of the inverse plasma-$\beta$, significant heating, but only moderate Lorentz factors. More specifically, in the HMF (LMF) case, the inverse plasma-$\beta$ can readily exceed values $\mathcal{O}(10)$ ($\mathcal{O}(5)$) within the funnel region, leading to a strong Parker instability, which transports highly magnetized material upwards. Furthermore, the Lorentz factors measured at about $300\,{\rm km}$ are only $W=1.02$ for the HMF scenario, to which correspond asymptotic values $W_{\infty}:=-hu_{t}=1.17$ that are much smaller than those expected in the phenomenology of short GRBs. Finally, the density contours reveal that the merger remnant expands in the vertical ($z$-axis) direction, with matter accumulating in this region. This accumulation contributes to the eventual weakening of the strong Poynting flux (see also Sec. V.3).
In order to gain a full 3D view of the propagation of these outflows, we have analyzed their properties via the projection of the Poynting flux on the two-sphere with coordinate radius $r\simeq 300\,{\rm km}$. In particular, we show in the top panel of Fig. 10 the projection at $\bar{t}=100\,\rm ms$, while the bottom panel refers to a much later time of $\bar{t}=170\,\rm ms$. Furthermore, for each panel, the left and right parts refer to the LMF and HMF scenario, respectively; note that because the corresponding EM luminosities are considerably different in the LMF and HMF, we have multiplied by a factor of ten the Poynting luminosity of the LMF to make it appear on the same colormap.
Overall, Fig. 10 allows us to appreciate that in the LMF case, the Poynting flux is not highly collimated at $\bar{t}=100\,\rm ms$, though it remains confined within a region with a polar angle of $\theta<\pi/3$. As the outflow evolves, it becomes increasingly collimated, with the Poynting flux being eventually confined within a narrower region of $\theta<\pi/6$ at the time of its peak value, which occurs at $\bar{t}=170\,\rm ms$ (see also Fig. 11). In the HMF scenario, on the other hand, the Poynting flux is consistently stronger, resulting from earlier breakup and a more extended distribution of magnetic-field structures, along with disk contributions (see also Sec. V.3) and strong flares from higher density regions (see also Sec. V.5). Interestingly, at $\bar{t}=170\,\rm{ms}$, the angular distributions of the Poynting flux in the polar region are remarkably similar in both the LMF and HMF cases. The primary distinction lies in the magnitude, which is larger by a factor of $\mathcal{O}(4)$ when strong initial magnetic fields are present. In contrast, at lower latitudes, multiple peaks associated with low-latitude flares are observed in the HMF case (see also Fig. 12). However, no such strong peaks appear in the LMF case, where the angular distribution of the Poynting flux remains comparatively flat. Interestingly, the rotation of the merger remnant and the flares that take place in its outer layers, are clearly imprinted in the fluctuations that are visible on the two-sphere projections. Hence, this phenomenology opens the possibility of learning about the rotation rate of the remnant by carefully studying the time statistics of these fluctuations in those (possibly rare) cases where they are measurable.
Figure 11: Evolution of the Poynting luminosity (top panel) and of the cumulative unbound mass (bottom panel) as computed on a two-sphere at a radius of $300\,\mathrm{km}$ for the LMF (red solid line) and the HMF (black solid line). Shown with dashed lines are the breakout times in the two scenarios, while the dot-dashed lines mark the times corresponding to the outflow properties shown in Fig. 10 and Fig. 12.
Figure 12: Angular distribution of the Poynting flux (top panel) and velocity distribution of the unbound matter (bottom panel) as measured by a detector at radius of $300\,\rm km$. Lines of different colours refer to either the LMF (red line) or to the HMF (black line) scenario. Furthermore, lines of different shadings show the distributions at different times, namely at $\bar{t}_{1}=100\,\rm ms$ (lighter lines) or at $\bar{t}_{2}=170\,\rm ms$ (darker lines), so as to show how these quantities evolve in time (see also Fig. 10 and Fig. 11).
[…]
Figure 13: Distributions of the Poynting fluxes (top row) and of the inverse plasma-$\beta$ (bottom row) as projected on two-sphere of four different radii (i.e., from left to right: $15\,\rm km$, $75\,\rm km$, $150\,\rm km$, and $300\,\rm km$) at $\bar{t}_{2}=170\,\rm ms$. The data refers to the LMF case; a similar evolution is shown in Fig. 14 for the HMF case.
All things considered, the differences between the EM luminosity in the HMF and LMF cases can be ascribed to three main factors. First, the earlier breakout of the magnetic field in the HMF case allows for the propagation of a larger amount of EM energy from the remnant, cleaning up the funnel; by contrast, in the LMF case the EM energy remains confined close to the remnant for a longer time. Second, in the HMF case, flares originate from a higher-density region of the magnetar, causing stronger eruptions than those produced in the LMF case, where flares are rooted in the relatively low-density region. Third, the Parker instability criterion is frequently met in the low-latitude regions of the HMF scenario (see bottom panel of Fig. 7), contributing to stronger Poynting flux; on the contrary, the criterion is marginally met and only in the funnel in the LMF case. Interestingly, a stronger Poynting flux produced by a stronger magnetic field was reported also by Ref. Bamber et al. (2025), though their simulations were conducted on a much shorter timescale.
A more quantitative measure of the anisotropy of the emission and of the degree of collimation of the outflows is presented in top panel of Fig. 12, which reports the polar distribution of the Poynting luminosity in the case of the LMF (red solid lines) and of the HMF (black solid lines) scenarios; furthermore, different shadings refer either to an early time at $\bar{t}_{1}=100\,\rm ms$ (lighter lines) or to a late one at $\bar{t}_{2}=170\,\rm ms$ (darker lines). Concentrating on the latter, it is clear that the degree of collimation is the same in the LMF and HMF cases, but that the intensity of the luminosity differs by an order of magnitude. Also quite evident when comparing the distributions at different times is that the degree of collimation increases with time, as the merger remnant reaches a quasi-stationary evolution and the fluctuations related to the ejection of matter have been washed out.
Also reported in Fig. 11, but in the lower panel, is the evolution of the ejected mass (see Appendix A for details on the definition and calculation). Clearly, a substantial amount of mass is ejected by the remnant over the timescale of the simulation and this amounts to a total of $M_{\rm ej}\sim 0.11\,M_{\odot}$ for the LMF case and $\sim 0.17\,M_{\odot}$ for the HMF case, thus indicating that strong magnetic fields in the merger remnant greatly facilitate the mass ejection. Less obvious is why for $20\lesssim\bar{t}\lesssim 50\,{\rm ms}$ the ejected mass is larger in the LMF case than in the HMF case. We believe that this is because of the larger braking action in the HMF case, that converts the kinetic-energy reservoir in the remnant into toroidal magnetic field. This braking is smaller for the LMF scenario and hence the kinetic energy is employed more efficiently to eject unbound matter. Part of the reason may stem from the pinch effect of the strong toroidal magnetic field (see, e.g.,  Frieben and Rezzolla (2012)), which has already reached very large values by $\bar{t}\sim 20\,\rm{ms}$ in the HMF case. This effect can limit matter expansion in the disk region, the primary source of the ejected mass.
Finally, shown in the bottom panel of Fig. 12 is the velocity distribution of the unbound mass with the same colour convention used for the distribution of the Poynting luminosity in the top panel. Note that for both magnetic-field configurations, the distribution reveals two components: a subdominant high-velocity component, i.e., with $0.1\lesssim v\lesssim 0.15$, associated with the matter ejected from the polar region, and a dominant low-velocity component, i.e., with $v\lesssim 0.1$, primarily contributed by material from the disk.
Figure 14: The same as in Fig. 13, but for the HMF case.
Note also that the late-time distribution of the HMF case has lost part of its high-velocity tail exhibited at $\bar{t}_{2}=170\,\mathrm{ms}$ and that reached values $v\lesssim 0.2$. This is because the most energetic emission in the case of large initial magnetic fields takes place early on and for $\bar{t}\lesssim 120\,\mathrm{ms}$. Interestingly, the opposite takes place for the LMF case, where the late-time distribution gains a high-velocity component. This is because the LMF Poynting flux reaches its peak value at $\bar{t}_{2}=170\,\mathrm{ms}$, as buoyant material driven by Parker instability clears the funnel region, allowing ejected matter to escape more rapidly. However, the corresponding Lorentz factors in all cases remain rather low and smaller that those reported in Refs. Kiuchi et al. (2024); Most and Quataert (2023); Musolino et al. (2024a), as the funnel region is heavily polluted by baryons, and neutrino emission is neglected in the simulation. Including the latter will boost the material to larger relativistic velocities with terminal Lorentz factors as large as $2-20$ Kiuchi et al. (2024); Musolino et al. (2024a), which however fall short of the values $\mathcal{O}(100)$ expected from the phenomenology of short GRBs Duncan and Thompson (1992); Usov (1992); Dai and Lu (1998a); Zhang and Mészáros (2001).
V.5 Dynamics of the flares near the HMNS
In the previous Sec. V.3, we demonstrated that flares are generated by the Parker instability and analyzed their formation sites. Here, we extend our analysis to their dynamics in the vicinity of the HMNS. As shown in Fig. 8, the Parker instability gives rise to magnetic-field arc structures, which are further inflated by convective motion. Due to the highly differential rotation of the remnant, the footpoints of these arcs undergo differential motion, leading to the formation of twisted magnetic loops. These loops rise as a result of magnetic buoyancy, transporting magnetically dominated material upward. This process heats the surrounding matter and drives a strong Poynting flux, introducing additional variability in the EM emission. Furthermore, these eruptions play a crucial role in shaping the structured poloidal magnetic field in the polar region (see Sec. V.4). This configuration facilitates mass outflow and angular momentum transport along the $z$-direction via the magnetocentrifugal mechanism.
To further quantify these dynamical processes, we present in Figs. 13 and 14 a series of projections on two-spheres in the LMF and HMF cases, respectively. More specifically, for each figure, in the top part we show the projections of the Poynting luminosity on different extraction radii, while in the bottom one we report the corresponding projections for the inverse plasma-$\beta$. We should note that capturing the evolution of these flares and of their highly dynamical propagation is inherently challenging; thus, our approach provides an indirect yet insightful perspective on the underlying dynamics. Notably, we find that eruptions near the HMNS significantly enhance the Poynting flux in both the LMF and HMF cases. This enhancement is evident as regions with large $\beta^{-1}$ consistently exhibit a higher Poynting flux than their surroundings.
At small radii, i.e., $r\sim 15\,\rm{km}$, the strong Poynting flux primarily originates from flares, namely, matter outflows carrying intense magnetic fields during the violent oscillation phase of the HMNS. At the eruption, magnetic energy in the flare is converted to kinetic energy, but as they propagate outward, they naturally interact with the ambient plasma, transferring part of their kinetic energy, but also expanding and cooling, and ultimately weakening their strength, as a result, the Poynting flux at large radii, i.e., $r\sim 300\,\rm{km}$, has been reduced significantly. Furthermore, we observe distinct morphological differences across the different two-spheres: in the polar regions, structures tend to form ring-like patterns, whereas at lower latitudes, they resemble belts. This distinction arises naturally due to the rapid rotation of the remnant and the large degree of anisotropy in the ejection of matter.
Obviously, together with this common qualitative phenomenology, the flares in the LMF and HMF scenarios measured on the two-spheres differ quantitatively in the frequency and strength becoming both more and stronger as the initial magnetic field is increased, as already encountered with Figs. 7 and 8. This is not surprising and indeed it is consistent with the expectation that flares simply reflect the breaking of a local hydromagnetic equilibrium in the outer layers of the merger remnant. As such, stronger magnetic fields naturally favour the breaking of the local hydromagnetic equilibrium.
In summary, these intermittent eruptions from the HMNS, that have been reported in other works and scenarios Siegel et al. (2014); Kiuchi et al. (2018); Most and Philippov (2020); Most and Quataert (2023); Most (2023); Carrasco et al. (2019); Nathanail et al. (2020, 2021b); Mahlmann et al. (2023); Musolino et al. (2024a), contribute to strengthen the Poynting flux, to the formation of a structured poloidal magnetic field in the polar region, and to aiding collimated outflow formation. These considerations hold true even if the magnetic fields are not very strong, but the frequency and strength of the flares is obviously closely related to the intensity of the magnetic energy. If the flares reported in this work were to be sustained over long periods of time and detected in the emission before and after a short (or even a long) GRB, they could reveal the presence of a highly magnetized, long-lived neutron star following a BNS merger (see also relevant discussion in Refs. Dai and Lu (1998b); Fan and Xu (2006); Gao and Fan (2006); Dai et al. (2006); Dall’Osso et al. (2011); Rowlinson et al. (2013); Wang et al. (2024)). In turn, this information could be used to further constrain the maximum neutron star mass and EOS and some intriguing observational evidence has possibly already emerged Chirenti et al. (2023); Yang et al. (2025). We note that, aside from the MHD effects identified by Refs. Siegel et al. (2013); Combi and Siegel (2023); Kiuchi et al. (2024); Most (2023); Musolino et al. (2024a), which suggest that the MRI and dynamo processes can drive relativistic outflows or flaring, our results reveal that the Parker instability can directly or indirectly contribute to flaring and relativistic outflows both before and after a short GRB. Clearly, future simulations containing full neutrino and EM radiative transfer process are needed to sharpen this picture and obtain more precise predictions on the role played by the Parker instability and by magnetic flares on the phenomenology of BNS mergers, but also of long GRBs.
VI Conclusions
Much of the future progress in the modelling of the merger of binary systems comprising at least a neutron star will depend very strongly on the ability to carry out simulations that model the system, either in the inspiral or after the merger, for timescales that are $\mathcal{O}(100-1000)$ times the dynamical one. Such timescales, however, are prohibitively expensive for the standard approaches in numerical relativity to simulate BNS and BHNS mergers.
Our approach to tackle these challenges consists in a hybrid approach that we have recently developed and that combines, via a hand-off transition, a fully general-relativistic code, i.e., FIL, with a more efficient code, i.e., BHAC+, making use of the conformally flat approximation (CFC). A crucial aspect of the CFC approach that makes it computationally attractive is that the corresponding field equations are elliptic and need not to be updated at every time-step over which the GRMHD equations are solved. Rather, they can be solved every $10$-$40$ time-steps, thus decreasing computational costs. In addition, when solving the set of GRMHD equations, the time-step size needs not be CFL-limited by the speed of light, but by the relevant MHD speeds, which are $\sim 30\%$ smaller. As a result, speed-ups of the order of $3-4$ are possible in our hybrid approach and, in turn, simulations on timescales $\mathcal{O}(100-1000)\,{\rm ms}$ can be carried out with much reduced computational costs.
While our hybrid approach was first presented and discussed in detail in Ref. Ng et al. (2024a), we have here reported important additional developments of BHAC+ consisting of the inclusion of gravitational-wave radiation-reaction contributions and of higher-order formulations of the equations of GRMHD. Both improvements have allowed us to explore the BNS merger remnants with high accuracy and over timescales that would have been computationally prohibitive with FIL.
More specifically, using these improvements, we have stress-tested the hybrid approach carrying out the hand-off procedure under different conditions of the spacetime dynamics and GW content. Our findings indicate that this approach yields highly accurate evolutions for the postmerger remnant, capturing both overall dynamics and magnetic field evolution effectively. The agreement between the two codes greatly improves as the GW amplitude decreases, especially when GW amplitude has decreased to $1/16$ of its peak value. We suggest the latter as the reference criterion for performing the hand-off as it is gauge-invariant, model independent, and only weakly dependent on resolution.
Besides reporting computational developments, we have applied the new hybrid approach to investigate the impact of the magnetic-field strength on the long-term and high-resolution evolutions of the “magnetar” resulting from the merger of two neutron stars with a realistic EOS. More specifically, using the TNTYST EOS, we have simulated magnetars formed after BNS mergers exploring two different magnetic-field strengths representing a low and a high-magnetic-field scenario (LMF and HMF, respectively) with a resolution of $150\,\rm m$ over a simulation period of $200\,\rm ms$. Interestingly, in the LMF case, the winding time is prolonged, and the magnetic field generated through winding is eventually larger than that of the HMF case. This behaviour can be attributed to a stronger back-reaction of the toroidal magnetic-field tension and hence to the suppression of small-scale turbulence in the high magnetic-field scenario. Furthermore, while in both scenarios we observe the weakening or suppression of differential rotation, the latter takes place on smaller timescales when the magnetic field is stronger, as expected.
Interestingly, we find that the physical conditions in the outer layers of the merger remnant close to the polar axis can lead to a global breakout of the plasma as a result of the development of the Parker instability (as also reported in Ref. Musolino et al. (2024a)). This breakout leads not only to an enhancement of the Poynting flux, but also to the occurrence of local eruptions of plasma, or “flares”, that are present in both the LMF and HMF scenario, although with different frequency and strengths. These flares, that are driven mostly by the Parker instability and have been reported in other works and scenarios Siegel et al. (2014); Kiuchi et al. (2018); Most and Philippov (2020); Most and Quataert (2023); Most (2023); Carrasco et al. (2019); Nathanail et al. (2020, 2021b); Mahlmann et al. (2023); Musolino et al. (2024a), contribute to strengthen the Poynting flux, to the formation of a structured poloidal magnetic field in the polar region, and to aiding collimated outflow formation. As a result, we measure Poynting luminosities reaching $1.9\times 10^{52}\,\rm erg~{}s^{-1}$ ($1.4\times 10^{51}\,\rm erg~{}s^{-1}$) and collimated within a half-opening angle of $\pi/6$ for the HMF (LMF) scenario. Independent of the strength of the initial magnetic field, the growth of the magnetic energy slows down and the main difference is in the time this takes place, which can be $\mathcal{O}(30)\,{\rm ms}$ for the HMF or $\mathcal{O}(100)\,{\rm ms}$ for the LMF. Hence, by the time the simulations are ended after $\sim 200\,{\rm ms}$ the collimated Poynting fluxes are weakened. The inclusion of neutrino transport, and the consequent clearing of the funnel operated by the neutrino-driven winds, are likely to counteract the decrease in Poynting flux and possibly increase the impact of the flares on the variability of the EM luminosity.
As a final remark we note that if these eruptions were to be sustained over long timescales and detected in the emission before and after a short GRB, they could reveal the presence of a long-lived magnetar in a BNS merger and potentially provide information of the properties of the remnant and its EOS. While these prospects are very exciting and some intriguing observational evidence may have already emerged Chirenti et al. (2023); Yang et al. (2025), future simulations are needed to confirm this picture. In particular, including a treatment of neutrino and EM radiation transfer and carrying out these simulations on even longer timescales, will certainly help attain a more realistic description of the long-term evolution of the remnant from BNS mergers. We plan to report on these improvements in future work.
Acknowledgements.
We thank R. Oechslin and G. Schaefer for the extensive exchange on the inclusion of the RR terms and E. Most and C. Musolino for useful discussions. This research is supported by the ERC Advanced Grant “JETSET: Launching, propagation and emission of relativistic jets from binary mergers and across mass scales” (grant No. 884631), by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the CRC-TR 211 “Strong-interaction matter under extreme conditions” – project number 315477589 – TRR 211, by the GSI Helmholtzzentrum für Schwerionenforschung, Darmstadt as part of the strategic R&D collaboration with Goethe University Frankfurt and by the State of Hesse within the Research Cluster ELEMENTS (Project ID 500/10.006). JLJ acknowledges partial support by the Alexander von Humboldt Foundation. MC acknowledges support from the NSF grants PHY-2110338, PHY-2409706, AST-2031744 and OAC-2004044 as well as the NASA TCAN Grant No. 80NSSC24K0100. LR acknowledges the Walter Greiner Gesellschaft zur Förderung der physikalischen Grundlagenforschung e.V. through the Carl W. Fueck Laureatus Chair. The simulations were performed on HPE Apollo HAWK at the High Performance Computing Center Stuttgart (HLRS) under the grant BNSMIC.
Software.
The software employed in this work comprises the following codes: BHAC Porth et al. (2017); Olivares et al. (2019); Ripperda et al. (2019), BHAC+ Ng et al. (2024a), FIL Most et al. (2019); Chabanov et al. (2023); Musolino and Rezzolla (2024); Chabanov and Rezzolla (2023); Ng et al. (2024), ETK Löffler et al. (2012), FUKA Papenfort et al. (2021). Furthermore, it makes use of the CompOSE (https://compose.obspm.fr) database for the handling of the EOS Typel et al. (2015).
Appendix A Diagnostic quantities
In what follows we provide some explicit expressions for the definition of quantities discussed extensively in the main text. Different implementations of the same formula may be used in FIL and BHAC+ because different gauge conditions and metric forms are adopted in these two codes.
We start with the baryon mass that is calculated as
$M_{\rm b}:=\int_{\mathcal{V}}W\rho\sqrt{\gamma}\,d^{3}\,x\,,$ (17)
where $W$ is the Lorentz factor of the fluid, and $\gamma$ is the determinant of the purely spacial metric. Similarly, the total internal energy is defined by
$E_{\rm int}:=\int_{\mathcal{V}}W\epsilon\rho\sqrt{\gamma}\,d^{3}\,x\,,$ (18)
where $\epsilon$ is specific internal energy. Another important energy considered in our analysis the total EM energy that in BHAC+ is defined as
$E_{\rm EM}:=\frac{1}{2}\int_{\mathcal{V}}(B_{i}B^{i}+E_{i}E^{i})\sqrt{\gamma}\,d^{3}\,x\,,$ (19)
where $B^{i}$ and $E^{i}$ are the magnetic and electric fields measured in the Eulerian frame, respectively. We note that the equivalent expression in FIL contains a multiplicative factor $1/4\pi$ which comes from the different EM units used in the two codes Most et al. (2019); Porth et al. (2017). In addition, the EM energy in the toroidal magnetic field is computed as
$E_{\rm EM,tor}:=\frac{1}{2}\int_{\mathcal{V}}(B_{\phi}B^{\phi}+E_{\phi}E^{\phi})\sqrt{\gamma}\,d^{3}\,x\,,$ (20)
while the poloidal EM energy is calculated as the difference between the total and toroidal EM energies. Finally, since we consider the ideal-MHD limit, the following relation holds
$E^{i}=\gamma^{-1/2}\eta^{ikl}v_{k}B_{l}\,,$ (21)
where $v_{k}$ is the Eulerian velocity, and $\eta^{ikl}$ is the three-dimensional Levi-Civita symbol defined as
$\eta^{ikl}=\eta_{ikl}:=\left\{\begin{array}[]{ccc}+1&\mathrm{if}&(i,k,l)\mathrm{~{}even~{}permutation}\,,\\ -1&\mathrm{if}&(i,k,l)\mathrm{~{}odd~{}permutation}\,,\\ 0&\mathrm{if}&\mathrm{else}\,.\end{array}\right.$ (25)
In terms of radiative quantities, the Poynting flux is calculated on a two-sphere of coordinate radius $r$ as
$L_{\rm Poyn}:=\oint_{dS}\left(b^{\mu}b_{\mu}u^{r}u_{0}-b^{r}b_{0}\right)\alpha\psi^{6}r^{2}\sin{\theta}d\phi d\theta\,,$ (26)
where $b^{\mu}$ is the magnetic field in the fluid frame and $u^{\alpha}$ is the four-velocity in the Eulerian frame. The unbound rest-mass is calculated in a similar way using
$\dot{M}_{\rm ej}=-\oint_{dS}W\tilde{\rho}\left(\alpha v^{r}-\beta^{r}\right)\psi^{6}r^{2}\sin{\theta}d\phi d\theta\,,$ (27)
[…]
Appendix B Details on the inclusion of RR terms
Although the first inclusion of RR terms was presented in Ref. Oechslin et al. (2007), a clear derivation of the PN corrections and their exact coupling to a GRMHD code with CFC approximation has not been presented yet. In addition, since unreported typos Oechslin appear in Eq. (A.37) of Ref. Oechslin et al. (2007), we report briefly the expressions for the PN-corrected components of the metric that are implemented in BHAC+ and refer the reader to Refs. Blanchet et al. (1990); Faye and Schäfer (2003) for the full derivations, with the caveat that these references derive quantities to be implemented within purely PN numerical codes rather than in codes that solve the equations of GRMHD or that adopt the CFC approximation.
The most important RR correction in our CFC approach is contained in Eq. (11) for the metric component $g_{00,{{}_{\rm RR}}}$ and to discuss how this is obtained we first need to recall the 2.5 PN order formalism with 3.5 PN corrections derived in Ref. Faye and Schäfer (2003) and which improves upon the formalism in Ref. Blanchet et al. (1990). For convenience, we will report explicitly the powers of $c$ in the PN expansion and follow the notation of Ref. Faye and Schäfer (2003) if the variables are not defined in this paper, where the metric $g_{\mu\nu}$ splits into an odd $(g_{\mu\nu})_{({\rm odd})}$ and an even $(g_{\mu\nu})_{({\rm even})}$ part depending the parity of the terms they generate in the equations of motion of PN hydrodynamics. For example, $(g_{00})_{(7)}$, $(g_{0i})_{(6)}$ or $(g_{ij})_{(5)}$ belong to the odd part, while $(g_{00})_{(2)}$, $(g_{0i})_{(3)}$ or $(g_{ij})_{(2)}$ to the even one; note that both the odd and even parts include 2.5 and 3.5 PN corrections (see Faye and Schäfer (2003) for more details).
We start therefore with considering the components of the RR part of the gravitational field at 2.5 PN and 3.5 PN orders: $(g_{00})_{(7),{{}_{\rm RR}}}$ and $(g_{00})_{(9),{{}_{\rm RR}}}$ [see Eqs. (4.10a) and (4.18a) in Faye and Schäfer (2003)]. Given the definitions of the PN potentials $\mathcal{U}_{*}$, $\mathcal{U}_{*i}$, and $\mathcal{R}$ in Eqs. (7a)–(7c), and considering only $Q^{[3]}_{ij}$ related terms, we follow Oechslin et al. (2007) and replace $D^{*}$ with the 1 PN + 3.5 PN mass density, i.e.,
$\sigma:=T^{00}+T^{ii}\,,$ (28)
and replace $v^{i}$ with the covariant momentum per unit rest-mass in the fluid frame
$w_{i}:=hu_{i}=v^{i}+\mathcal{O}(1/c^{2})\,,$ (29)
where it includes the 1 PN correction of $v^{i}$.
Applying these replacements and keeping leading order terms with dependency of $Q^{[3]}_{ij}$, the expression for the $(g_{00})_{(7),{{}_{\rm RR}}}$ metric component can be simplified as follows
$\left(g_{00}\right)_{(7),{{}_{\rm RR}}}$ $=\frac{4}{5c^{7}}\left(-Q_{kl}^{[3]}x^{k}\partial_{l}\mathcal{U}_{*}+\int\frac{d^{3}\mathbf{y}}{|\mathbf{x}-\mathbf{y}|}Q_{kl}^{[3]}x^{k}\partial_{l}\sigma\right)$ (30)
$=\frac{4}{5c^{7}}\left(-Q_{kl}^{[3]}x^{k}\partial_{l}\mathcal{U}_{*}+\mathcal{R}\right)\,,$
where the Poisson equation for $\mathcal{R}$ is given in Eq. (7c).
Similarly, the expression for the metric correction $(g_{00})_{(9),\mathrm{RR}}$ can be written as
$\left(g_{00}\right)_{(9),{{}_{\rm RR}}}=$ $\frac{4}{5c^{9}}\left\{-I_{2kl}^{[3]}x^{k}\partial_{l}\mathcal{U}_{*}+Q_{kl}^{[3]}x^{k}\left(-\partial_{l}\mathcal{U}_{2}+2\mathcal{U}_{*}\partial_{l}\mathcal{U}_{*}\right)+Q_{kl}^{[4]}x^{k}\left(-\frac{1}{2}x^{l}\partial_{l}\mathcal{U}_{*}+A_{*l}\right)+\frac{5}{126}Q_{klm}^{[5]}x^{k}x^{l}\partial_{m}\mathcal{U}_{*}\right.$ (31)
$+Q_{kl}^{[5]}x^{k}\left(\frac{17}{42}x^{l}x^{m}\partial_{m}\mathcal{U}_{*}-\frac{11}{42}r^{2}\partial_{l}\mathcal{U}_{*}\right)-\frac{8}{9}\epsilon_{klm}S_{mn}^{[4]}x^{l}x^{n}\partial_{k}\mathcal{U}_{*}-2\mathcal{U}_{*}\int\frac{d^{3}\mathbf{y}}{|\mathbf{x}-\mathbf{y}|}Q_{kl}^{[3]}y^{k}\partial_{l}\sigma$
$+\int\frac{d^{3}\mathbf{y}}{|\mathbf{x}-\mathbf{y}|}\left[I_{2kl}^{[3]}y^{k}\partial_{l}\sigma+Q_{kl}^{[3]}y^{k}\left[\sigma\partial_{l}\mathcal{U}_{*}+\partial_{l}\left(\sigma\delta\right)\right]-3\sigma w_{k}w_{l}Q_{kl}^{[3]}-\frac{5}{126}Q_{klm}^{[5]}y^{k}y^{l}\partial_{m}\sigma\right.$
$+Q_{kl}^{[4]}y^{k}\left(\frac{1}{2}y^{l}\partial_{t}\sigma-4\sigma w_{l}\right)+Q_{kl}^{[5]}y^{k}\left(-\frac{17}{42}y^{l}y^{m}\partial_{m}\sigma+\frac{11}{42}|\mathbf{y}|^{2}\partial_{l}\sigma-\sigma y^{l}\right)-\frac{8}{9}\epsilon_{klm}S_{mn}^{[4]}y^{l}y^{n}\partial_{k}\sigma$
$\left.\left.-\sigma\int\frac{d^{3}\mathbf{y}^{\prime}}{\left|\mathbf{y}-\mathbf{y}^{\prime}\right|}\left(Q_{kl}^{[3]}y^{k}\partial_{l}\sigma\right)\left[\mathbf{y}^{\prime}\right]\right]\right\}$
$\approx$ $\frac{4}{5c^{9}}\left(2Q_{kl}^{[3]}x^{k}\mathcal{U}_{*}\partial_{l}\mathcal{U}_{*}-2\mathcal{U}_{*}\mathcal{R}+\mathcal{R}_{2}\right)\,,$
where, after grouping the terms inside the second integral of $\int d^{3}\mathbf{y}$, a Poisson equation for the potential $\mathcal{R}_{2}$ is expressed as [see Eq. (4.31cc) in Ref. Faye and Schäfer (2003)]
$\Delta\mathcal{R}_{2}=$ $-4\pi\left[I_{2}{}_{kl}^{[3]}x^{k}\partial_{l}\sigma+Q_{kl}^{[3]}x^{k}\left[\sigma\partial_{l}\mathcal{U}_{*}+\partial_{l}\left(\sigma\delta\right)\right]\right.-3\sigma w_{k}w_{l}Q_{kl}^{[3]}-\frac{5}{126}Q_{klm}^{[5]}x^{k}x^{l}\partial_{m}\sigma$ (32)
$+Q_{kl}^{[4]}x^{k}\left(\frac{1}{2}x^{l}\partial_{t}\sigma-4\sigma w_{l}\right)+Q_{kl}^{[5]}x^{k}\left(-\frac{17}{42}x^{l}x^{m}\partial_{m}\sigma+\frac{11}{42}r^{2}\partial_{l}\sigma-\sigma x^{l}\right)$
$\left.-\frac{8}{9}\epsilon_{klm}S_{mn}^{[4]}x^{l}x^{n}\partial_{k}\sigma-\sigma\mathcal{R}\right]$
$\approx$ $-4\pi\sigma\left[Q_{kl}^{[3]}x^{k}\partial_{l}\mathcal{U}_{*}-3w_{k}w_{l}Q_{kl}^{[3]}-\mathcal{R}\right]\,.$
We show that the corresponding expression of Eq. (4.31cc) in Ref. Faye and Schäfer (2003) contains a typo in the sign of the term $3\sigma w_{k}w_{l}Q_{kl}^{[3]}$, as it is easy to deduce when starting from Eq. (4.18a) in Ref. Faye and Schäfer (2003).
When focusing solely on the leading-order contributions involving $Q^{[3]}_{ij}$, namely, $\left(g_{00}\right)_{(7),\mathrm{RR}}$ and $\left(g_{00}\right)_{(9),\mathrm{RR}}$, the corresponding corrections to the RR metric component, $g_{00,\mathrm{RR}}$, can be expressed as:
$g_{00,{{}_{\rm RR}}}$ $\approx\left(g_{00}\right)_{(7),\mathrm{RR}}+\left(g_{00}\right)_{(9),\mathrm{RR}}$ (33)
$=\frac{4}{5}\left[\left(-Q_{kl}^{[3]}x^{k}\partial_{l}\mathcal{U}_{*}+\mathcal{R}\right)+\left(Q_{kl}^{[3]}x^{k}\mathcal{U}_{*}\partial_{l}\mathcal{U}_{*}-2\mathcal{U}_{*}\mathcal{R}+\mathcal{R}_{2}\right)\right]$
$=-\frac{4}{5}\left[(1-2\mathcal{U}_{*})\left(Q_{kl}^{[3]}x^{k}\partial_{l}\mathcal{U}_{*}-\mathcal{R}\right)\right]+\frac{4}{5}\mathcal{R}_{2}\,,$
where we note that the second term, i.e., $(4/5)\mathcal{R}_{2}$, is different from the term reported as $-(8/5)\mathcal{U}_{7}$ in Eq. (A.37) of Ref. Oechslin et al. (2007), although our definition of $\mathcal{R}_{2}$ is the same as that of $\mathcal{U}_{7}$ in Ref. Oechslin et al. (2007) with an extra term $-3Q_{ij}^{[3]}w_{i}w_{j}$.
Finally, when computing the GW emission, it is possible to express the radiating parts of the metric by computing the Newtonian mass quadrupole formula for the matter distribution and express the second time derivative of the mass quadrupole $Q_{ij}$ analytically after exploiting the conservation of rest-mass as Blanchet et al. (1990)
$Q_{ij}^{[2]}=\Biggl{[}2\int d^{3}xD^{*}\left(V^{i}V^{j}+x^{i}\partial_{j}\mathcal{U}\right)\Biggr{]}^{\mathrm{STF}}+\mathcal{O}\left(\frac{1}{c^{2}}\right)\,.$ (34)
The corresponding GW strain at the pole in the two polarisations $+$ and $\times$ and at a distance $r$ is given by Ref. Dimmelmeier et al. (2005); Yip et al. (2023)
$h_{+}=$ $\frac{(Q^{[2]}_{xx}-Q^{[2]}_{yy})}{r}\,,$ (35)
$h_{\times}=$ $\frac{2Q^{[2]}_{xy}}{r}\,.$ (36)
Appendix C On the frequency of the spacetime update
As mentioned in Sec. II.1, one of the most significant advantages of the hand-off to BHAC+ is given by the considerably higher efficiency with which the evolution can be carried out. Among the various sources of this efficiency – which include: a better use of the memory and AMR, a reduced set of equations for the solution of the field variables, and a CFL constraint set by the MHD speeds rather than the speed of light – the most significant one is probably to be found in the frequency of the spacetime update. Indeed, in contrast with what happens in FIL (but also in all other full-GR MHD codes), the solution of the field equations needs not be made at every time-level at which the GRMHD equations are solved. We measure this speed-up in terms of the ratio $\chi:=\Delta t_{\rm met}/\Delta t_{\rm MHD}$, where $\Delta t_{\rm met}$ ($\Delta t_{\rm MHD}$) is the time interval between two successive metric (MHD) updates, and $\chi=1$ corresponds to the case in which the field equations are solved at every time-step. We normally take $\Delta t_{\rm met}\geq\Delta t_{\rm MHD}$ so that $\chi\geq 1$.
Figure 15 shows a comparison in the evolution of the maximum rest-mass density (left panel) and of the total EM energy (right panel) between FIL and BHAC+ for the hand-off case PM-D (see also Fig. 2). It is quite remarkable that the BHAC+ solutions both in terms of rest-mass density and of EM energy are only weakly dependent on the value of $\chi$ despite that a larger value leads to a proportionally smaller computational cost. More precisely, Tab. 3 shows the computational costs of two simulations carried out over one millisecond and having $\chi=1$ ($\Delta t_{\rm met}=5.0\times 10^{-2}~{}M_{\odot}$) and $\chi=40$ ($\Delta t_{\rm met}=2.0~{}M_{\odot}$) differ by a factor $1595/944\simeq 1.7$; this speedup becomes of a factor $3117/944\simeq 3.3$ for $\chi=40$ when comparing with FIL. These speedups cannot be attributed to the different AMR approaches in the two codes, i.e., box-in-box in FIL vs block-based quadtree-octree AMR in BHAC+, as the two codes have very similar running times when simulating spherical stars in the Cowling approximation (hence without the solution of the field equations) and with AMR grid structures that have the same finest resolution and comparable number of cells.
Tests $\Delta t_{\rm met}$ ${\rm Cost~{}for~{}1~{}ms}$
$[M_{\odot}]$ $[\rm{CPU~{}hr}~{}\rm{ms^{-1}}]$
$\chi=1~{}~{}$ (FIL) $1.3\times 10^{-2}$ $3117$
$\chi=1~{}~{}$ (BHAC+) $5.0\times 10^{-2}$ $1595$
$\chi=2~{}~{}$ (BHAC+) $1.0\times 10^{-1}$ $1222$
$\chi=10$  (BHAC+) $5.0\times 10^{-1}$ $1025$
$\chi=40$  (BHAC+) $2.0$ 944944~{}944944
$\chi=\infty$ (BHAC+/FIL) $\infty$ $937$/$982$
Table 3: CPU hours required per millisecond of simulation time in BHAC+ and the corresponding size of the metric-update time-step for different values of $\chi$, as measured with on the PM-D hand-off. Also reported are the corresponding values for FIL, for which $\chi=1$. Note that all simulations have the same resolution with the finest grid spacing of $300\,\rm{m}$.
Figure 15: Evolution of the maximum rest-mass density (left panel) and of the total EM energy (right panel) as computed by FIL (red lines) and by BHAC+ with different metric update frequencies $\chi\geq 1$ (blue lines of different shade). The case $\chi=1$ ($\chi=40$) refers to when the metric update is done at every ($40$) evolution steps of the MHD equations. The data refers to the hand-off case PM-D and a CFL coefficient $\mathcal{C}_{\rm CFL}=0.2$ has been fixed. The green lines show the results relative to a fixed spacetime, i.e., $\chi=\infty$.
We should remark that the use of increasingly larger values of $\chi$ has also the effect of making the evolution slightly more dissipative, as can be appreciated by considering the amplitude of the post-hand-off oscillations, which are gradually decreased by the use of larger values of $\chi$. At the same time, the differences between the $\chi=1$ and $\chi=40$ runs are very small. Indeed, when comparing with the evolution of FIL, the $\chi=1$ and $\chi=40$ runs have averaged relative differences of $-0.3\%$ ($-3.7\%$) and $-0.2\%$ ($-4.0\%$) in the evolution of the maximum rest-mass density (total EM energy), respectively.
Finally, we note that the computational costs do not reduce significantly when increasing above $\chi\simeq 40$. Indeed, when omitting altogether the spacetime update, which corresponds to $\chi=\infty$ (Cowling approximation), the computational costs are decreased by an additional $\simeq 0.7\%$. However, as shown in Fig. 15, the evolution with a fixed spacetime is extremely inaccurate and is characterized by much larger central rest-mass densities and a more rapid decay of the EM energy due to the more efficient braking of the differential rotation, which causes the suppression of magnetic winding.
Figure 16: The same as in Fig. 4, but for the PM-C case, i.e., 2D comparison on the principal planes between the results of FIL and BHAC+ as computed after $20\,\rm ms$ from the hand-off of case PM-C.
In view of these tests and considerations, our default implementation is to use a rather small value of $\chi\simeq 10$ ($\Delta t_{\rm met}=0.5\,M_{\odot}$) within $\sim 20~{}\rm{ms}$ after hand-off and to increase it to $\chi\simeq 20-40$ to maximise accuracy and efficiency in the late-time evolution, when the merger remnant has reached a quasi-stationary configuration.
Appendix D Impact on the evolution with an early hand-off
In Fig. 16, we compare the results obtained using two codes, FIL and BHAC+, after $20\,\rm ms$ of the hand-off process for an earlier hand-off case PM-C comparing with the main text. Although the GW amplitude remains higher in this case compared to PM-D (as shown in Fig. 4), the differences between the two codes are quite similar. More specifically, we observe that key structures, such as the shell structure in temperature, the ring structure of the Lorentz factor on the equatorial plane, the wing structure of the Lorentz factor on the $(x,z)$-plane, and the strong toroidal magnetic field exhibit a good match between the two codes. Moreover, these prominent structures do not change significantly relative to the PM-D case, suggesting that the merger remnant is already well stratified. The poloidal magnetic field appears slightly more turbulent during this period in both codes, leading to a less precise match both in terms of magnetic-field topology and of the EM energy (see Fig. 2). Overall, Fig. 16 reveals that an early hand-off is possible and does not lead to dramatic changes in the dynamics of the postmerger remnant, especially over timescales $\mathcal{O}(1)\,{\rm s}$. At the same time, it provides guidance in preferring, whenever possible, a hand-off that is performed $\sim 30\,\rm ms$ after merger.
Appendix E On the conformal flatness of BNS postmerger spacetimes
The extensive and detailed comparisons between the evolutions from the full-GR code FIL and the xCFC code BHAC+ discussed in this paper have provided ample evidence that the CFC approximation provides an accurate description of the spacetime evolution of the postmerger object. However, it remains an interesting question to assess and quantify the degree of “conformal flatness” of the spacetime generated by the remnant. To this scope, we recall that the Cotton-York (CY) tensor $\bm{C}$ is a rank-2 tensor which consist of third-order derivatives of the spatial metric $\mathbf{\gamma}$ and is constructed from the more general Cotton tensor $\bm{\mathcal{C}}$, which, instead, is of rank three. Its definition is given by Gourgoulhon (2007)
${C}^{ij}:=$ $-\frac{1}{2}\frac{1}{\sqrt{\gamma}}\eta^{ikl}\mathcal{C}_{mkl}\gamma^{mj}$ (37)
$=$ $\frac{1}{\sqrt{\gamma}}\eta^{ikl}D_{k}\left(R^{j}_{\phantom{j}l}-\frac{1}{4}R\delta^{j}_{\phantom{j}l}\right)\,,$ (38)
where the operator $D_{k}$ denotes the covariant derivative with respect to the spatial metric. The tensors $R^{j}_{\phantom{j}l}$ and $R$ are the Ricci tensor and scalar, respectively, with respect to the spatial metric.
One of the most important properties of the CY tensor is directly inherited from the Cotton tensor, namely, as the vanishing of the Cotton tensor $\bm{\mathcal{C}}$ is a necessary and sufficient condition for the metric to be conformally flat Gourgoulhon (2007), conformally flat spacetimes admit also a vanishing CY tensor. Hence, following similar approaches as in Refs. Miller et al. (2004); Ott et al. (2007); Iosif and Stergioulas (2014) we will make use of this property in order to quantitatively assess how far the spacetime of merging BNS deviates from being conformally flat.
More specifically, we define the matrix norm of the CY tensor, $|H_{ij}|$, as the square root of the largest eigenvalue of $C_{ik}C^{k}_{\phantom{j}j}$ (see Ref. Iosif and Stergioulas (2014) for more details). In addition, $|H_{ij}|$ is normalised by the covariant derivatives of $R_{ij}$ in order to provide a local measure of the deviations from conformal flatness which leads us to the definition
$H:=\frac{|H_{ij}|}{\sqrt{D_{i}R_{jk}D^{i}R^{jk}}}\,.$ (39)
We note that in Ref. Iosif and Stergioulas (2014) it was observed that this normalization is not suitable for low-density matter distributions because the denominator $\sqrt{D_{i}R_{jk}D^{i}R^{jk}}$ can exhibit large variations; fortunately, this drawback has not emerged in our analysis. Finally, we construct a density-weighted integral using $H$ in order to obtain an integrated measure over the whole spatial slice $\Sigma_{t}$
$\left\langle H\right\rangle_{\rho}:=\frac{\int_{\mathcal{V}}\mathrm{d}^{3}x\sqrt{\gamma}\rho WH}{\int_{\mathcal{V}}\mathrm{d}^{3}x\sqrt{\gamma}\rho W}\,,$ (40)
such that large/small values of $H$ (and of its time derivative) should be interpreted as referring to spatial regions with large/small deviations from conformal flatness.
Figure 17: Evolution of various expressions of the normalised measure of the deviation from conformal flatness $H$ (left panel; see Eq. (39)) and of its time derivative (right panel). The black lines display the rest-mass density average using Eq. (40), while the blue lines display the maximum of $H$ over the entire domain. The red lines visualise the maximum value taken over a box centered around $x=y=0$ and having a side length of $\sim 2\times 33~{}\mathrm{km}$. Shown with vertical blue lines of different shades are the different hand-off times.
[…]
Furthermore, and as expected, during the merger all three measures increase rapidly. The quantity $\langle H\rangle_{\rho}$ achieves moderate values of $\sim 10^{-1}$, which agree with strongly differentially rotating relativistic stars Iosif and Stergioulas (2014). More specifically, these values correspond broadly to the configurations with the highest values of $T/|W|$ in Ref. Iosif and Stergioulas (2014), where $T/|W|$ denotes the ratio of the rotational over the gravitational binding energy; with such values of $T/|W|$, we expect the remnant to still be stable with respect to the dynamical $\ell=2,m=2$ barmode instability Shibata and Sekiguchi (2005); Baiotti et al. (2007); Manca et al. (2007), even in the presence of strong magnetic fields Franci et al. (2013). Between merger and the time PM-A, all three measures display strong and rapid variations as a result of the violent collision-and-bounce cycles of the two stellar cores. After PM-A, the quantity $\left\langle H\right\rangle_{\rho}$ remains almost constant as a result of weak GW emission, the dominant mechanism to change $T/|W|$.
We find that even more informative than the various measurements of $H$ presented in the left panel of Fig. 17 are the corresponding time-derivatives, which are shown instead on the right panel. In this case, it is straightforward to realise that those stages of the process where the spacetime is curved but not highly dynamical, i.e., during the inspiral and at late times in the postmerger, the time derivatives of $(H)_{\rm max,~{}33~{}km}$, and $\langle H\rangle_{\rho}$ are both very small and either slowly increasing before merger or slowly decaying after merger. These are indeed the stages when the GW emission is not very strong. As the binary separation decreases and the merger takes place, the time derivatives of both proxies follow a behaviour that is very similar to that associated with the gravitational radiation. These considerations also highlight a conclusion we had drawn also before, namely, that a late hand-off at about $30\,{\rm ms}$ after merger is preferred as the GW content in the spacetime is very small. At the same time, the CFC approach employed with an hand-off at about $20\,{\rm ms}$ after merger already presents a very good approximation to the full-GR spacetime.